### v1

# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

Abstract
With the rapid growth of multi-modal data from social media, short video platforms, and e-commerce, content-based retrieval has become essential for efficiently searching and utilizing heterogeneous information. Over time, retrieval techniques have evolved from Unimodal Retrieval (UR) to Cross-modal Retrieval (CR) and, more recently, to Composed Multi-modal Retrieval (CMR). CMR enables users to retrieve images or videos by integrating a reference visual input with textual modifications, enhancing search flexibility and precision. This paper provides a comprehensive review of CMR, covering its fundamental challenges, technical advancements, and categorization into supervised, zero-shot, and semi-supervised learning paradigms. We discuss key research directions, including data augmentation, model architecture, and loss optimization in supervised CMR, as well as transformation frameworks and external knowledge integration in zero-shot CMR. Additionally, we highlight the application potential of CMR in composed image retrieval, video retrieval, and person retrieval, which have significant implications for e-commerce, online search, and public security. Given its ability to refine and personalize search experiences, CMR is poised to become a pivotal technology in next-generation retrieval systems. A curated list of related works and resources is available at: Awesome Composed Multi-modal Retrieval.
Index Terms:
Composed Multi-modal Retrieval (CMR), Cross-modal Retrieval, Multi-modal Learning, and Vision-Language Models
1 Introduction
Nowadays, the unprecedented prosperity of social media, short video platforms, and e-commerce (such as TikTok, Facebook, Weibo, YouTube, Amazon, and Alibaba) has greatly promoted the rapid growth of multi-modal data, encompassing various modalities such as texts, images, and videos. Faced with this data flood composed of heterogeneous information, content-based retrieval, as an information processing technology to search and utilize these vast resources, not only plays a core role in the field of e-commerce, such as achieving accurate matching of goods, but also greatly enriches the user experience of the social media platform, allowing users to quickly and accurately find the content of interest. Due to its widespread applications in everyday life that almost impact everyone, content-based retrieval has attracted great attention from both academia and industry, whose research significance is self-evident.
Generally, the evolution of content-based retrieval technology has witnessed the transformation from Unimodal Retrieval (UR) to Cross-modal Retrieval (CR), and then to Composed Multi-modal Retrieval (CMR). Compared with the early-stage unimodal retrieval, which was limited to querying information within the same modality, e.g., image search for images, cross-modal retrieval has achieved remarkable accuracy and widespread application in the present era. This enables the search for semantically relevant content in one modality based on the instance query from another modality, e.g., using text search images or videos, allowing users to make full use of these heterogeneous data. In recent years, composed multi-modal retrieval has emerged as a thriving content-based retrieval technology. Within this technical framework, the system aims to discover images/videos that not only bear resemblance to the given reference image/video but also allow for specific modifications based on the provided textual feedback from the user. In this sense, CMR pioneers an advanced level of interactive and conditional retrieval mechanism, leveraging deep integration of visual and linguistic information. This integration greatly enhances the flexibility and precision of user-expressed search intents, injecting new vitality into domains such as internet search and e-commerce. Consequently, CMR exhibits vast potential and far-reaching impact as the next-generation content-based retrieval engine in real-world application scenarios.
A core of CMR is that it requires the synergistic understanding and composing of both input vision and language information as the multi-modal query. The earliest closely related studies of CMR are in the field of attribute-based fashion image retrieval, where the key difference is that the textual feedback in attribute-based fashion image retrieval is limited to a single attribute (e.g., ’mini’, ’white’, ’red’), while CMR is the natural language with multiple words (e.g., ’showing this animal of the input image facing the camera under sunlight’), which is more flexible yet challenging compared to the predefined set of attribute values. The pioneering CMR works are proposed in, where the input query is specified in the form of an image plus some natural language that describes desired modifications to the input image, and led a series of subsequent approaches. Current research in CMR is primarily focused on three main patterns: (1) supervised learning-based CMR (SL-CMR), which focuses on how to design a better combination mechanism of vision and language through supervised training of annotated data, (2) zero-shot learning-based CMR (ZSL-CMR), which focuses on how to simulate and build a visual-linguistic multi-modal information combination framework without annotated data, and (3) semi-supervised learning-based CMR (SSL-CMR), which focuses on how to enhance the learning of visual-linguistic combination via automatic generated data under a semi-supervised manner.
For the combined multi-modal retrieval models based on supervised learning, a notable characteristic is the requirement of annotated triplet data $(I^{r},T^{r},I^{t})$, where $I^{r}$ denotes the reference query image, $T^{r}$ denotes the relatively modified text, and $I^{t}$ denotes the ground truth target image. For the given inputs of $I^{r}$ and $T^{r}$, the typical pipeline of an SL-CMR model involves mining what content should be modified in the reference image $I^{r}$ according to the caption $T^{r}$, so as to learn a vison-language compositional embedding that encodes the information required to find the interested target image $T^{r}$. Thus, the main challenges faced by such approaches lie in addressing two issues: ”Where to see?”, which refers to attending to the content in the reference image that needs change, and ”How to change?”, which aims to modify the reference image based on the textual information while preserving the remaining information. In recent years, research on SL-CMR has primarily focused on three aspects: (1) data augmentation, focusing on alleviating the difficulty of triplet labeling and the defect of labeling noise; (2) model architecture, focusing on designing a better vision-language combiner via cross-modal feature alignment and fusion strategies, as well as the design of other novel frameworks that can be plugged; (3) loss optimization, focusing on the design of more reasonable feature combination constraints. Although supervised training relying on these carefully labeled data often offers high performance, SL-CMR inherently faces two shortcomings: 1) annotating such triplets is both difficult and labor-intensive, and 2) the supervised approaches trained on the collected limited and specific triplets are also hard for generalization.
Recently, composed multi-modal retrieval based on zero-shot learning has been proposed to address the above limitations. During the training process, the model is trained solely on easily obtainable image-text pairs, without the need for annotated triplets. Its training process usually revolves around learning the modality transformers, which simulate the combination of visual and linguistic information in testing. As a result, the training and testing phases typically involve different network structures. Thus, a main challenge of ZSL-CMR lies in designing transformation frameworks that are equivalent to supervised learning in the absence of supervision signals, aiming to maximize the zero-shot generalization ability as much as possible. To address this challenge, the academic community has developed strategies across three key aspects: (1) Image-side transformation: this approach focuses on learning the implicit or explicit visual-to-linguistic transformation using images as input. During testing, it converts the reference image into a query that can be integrated with the relative textual information; (2) Text-side transformation: in this approach, text is used as input to simulate image features, constructing a training framework that relies solely on language. During testing, the model directly takes image inputs; (3) External knowledge assistance: this approach explores additional knowledge to enhance details, such as attributes and colors, thereby improving retrieval performance. Despite these efforts, there is currently no comprehensive review that systematically organizes the existing research, challenges, and applications of this emerging field.
Although zero-shot combined multimodal retrieval does not rely on labeled data, its performance is often lower than supervised training, which brings obstacles to the application of the model. To alleviate this problem, some works have proposed a semi-supervised combined multimodal learning paradigm based on automatically generated triple data. In this setting, relying on the relatively easy-to-obtain image-text data, existing SSL-CMR work mainly generates triplet data from two aspects: (1) generating images, such as editing the input reference image according to the conditional text to create the target image; (2) generating text, such as describing the difference caption between the two input images. By generating these triplet data, the model can not only capture the combination of vision and language more accurately during learning, but also avoid the limitations of cumbersome annotation. Although the generated data may be noisy, it also combines the advantages of supervision and zero-shot learning, which is a promising direction.
Despite the rapid advancements in composed multi-modal retrieval, numerous challenges remain to be addressed. Issues such as fine-grained vision-language composition, retrieval efficiency in large-scale datasets, and generalization to unseen data require further exploration. Moreover, as real-world applications continue to expand, integrating CMR with cutting-edge technologies such as foundation models and generative AI holds great promise for enhancing retrieval performance. In this paper, we provide a comprehensive review of the existing research in CMR, summarize key challenges, and discuss potential future directions. We hope this work serves as a valuable reference for researchers and practitioners in the field, fostering further innovation in next-generation multi-modal retrieval systems.
2 Overview
In this section, we provide a concise overview of the taxonomy of current Composed Multi-modal Retrieval (CMR) methods. To better track the latest advancements in CMR, this survey categorizes existing methods from both technical and application perspectives. (1) From a Technical Perspective: we summarize different paradigms of multimodal compositional learning, including sophisticated module designs and innovative solutions. It covers the state-of-the-art techniques that address the challenges of understanding and integrating information from multiple modalities. (2) From an Application Perspective: we examine various application scenarios where multimodal compositional retrieval is employed, such as fashion e-commerce, video platforms, geospatial data, and others. It highlights how these methods are applied in real-world contexts, showcasing their versatility and practical utility. By providing a comprehensive summary and reference for researchers in both the technical and application aspects, this survey aims to facilitate deeper thinking of this field and promote further development in multimodal compositional retrieval. The insights gained from this survey can guide future research and innovation, helping to overcome the inherent complexities and challenges in this rapidly evolving area.
From a technical perspective, current CMR methods can be grouped into three main paradigms: supervised learning, zero-shot learning, and semi-supervised learning. Each paradigm addresses the integration of multiple modalities for effective retrieval, with distinct strategies and trade-offs.
Supervised Learning (SL-CMR) relies on annotated triplet data, such as (reference image $V_{ref}$, modification text $T_{mod}$, target image $V_{tar}$), to learn how to combine input modalities so that the combined content is semantically aligned with the target image. This process can be summarized as:
$X_{combined}=Combiner(V_{ref},T_{mod}),$ (1)
[…]
where $Combiner(\cdot)$ denotes the combination of visual and textual inputs, and $Supervisor(\cdot)$ enforces semantic alignment between the combined content and the visual target Supervised learning benefits from explicit supervision, enabling precise modality combination and high performance. However, it requires large-scale, labor-intensive annotation, which can limit its generalization to new domains.
Zero-Shot Learning (ZSL-CMR) overcomes the need for annotated triplets by using readily available image-text pairs to simulate modality combinations. The model learns to transform one modality into another, e.g., mapping a reference image to a textual representation:
$T_{ref}^{*}=Converter(V_{ref}),$ (2)
and then combines this with the modification text to form a composite query. This shifts the task to cross-modal text-to-image retrieval. While ZSL-CMR offers easy data acquisition and better generalization, it lacks direct supervision for modality combination, leading to lower performance on specific tasks compared to supervised learning.
Semi-Supervised Learning (SSL-CMR) combines elements of both supervised and zero-shot approaches. It uses modality generation (e.g., image-to-text or text-to-image) to automatically create triplet data from easily accessible image-text pairs $(V,T)$:
$\hat{V}_{ref},\hat{T}_{mod},\hat{V}_{tar}=Generator(V,T).$ (3)
The model then learns to combine visual and linguistic information, accounting for any noise in the generated data. Semi-supervised learning achieves high performance through automated data construction but faces challenges in generating large volumes of high-quality triplets, especially those with subtle semantic differences. Existing work highlights that generated content often contains errors and noise, which remains an ongoing issue.
Next, we will introduce multimodal combined retrieval based on supervised, zero-shot, and semi-supervised learning in Sections 3, 4, and 5, respectively. Then, in Section 6, we will introduce the applications of combined multimodal retrieval in different scenarios, including related datasets, evaluation metrics, and method performance. Finally, in Section 7, we conclude the paper and look forward to future research directions.
3 Supervised Learning-based CMR
Existing supervised learning-based methods for CMR can be categorized into three key dimensions: data, model architecture, and optimization objectives. First, as discussed in Section 2, the construction and annotation of multimodal triplet data pose significant challenges. Consequently, research in this area primarily focuses on automating data augmentation and addressing annotation noise. The goal is to develop techniques that can efficiently generate large-scale triplet datasets while ensuring the quality and reliability of annotations. Second, in terms of model architecture, existing studies investigate various strategies for integrating visual and linguistic modalities. These efforts emphasize the precise modeling and fusion of semantic information, focusing on retaining, modifying, and supplementing semantics during integration. The goal is to design models that effectively align and fuse both modalities, ensuring that the combined representation captures the intended modifications and preserves the relevant context. Finally, the optimization objectives under the triplet data framework represent another critical aspect. Building on contrastive learning, recent advancements have introduced improvements to objective functions, aiming to enhance the alignment between the fused input and the visual target, thereby improving retrieval performance. In the following, we provide a detailed review of the progress made in these areas, highlighting key contributions and ongoing challenges.
3.1 Data Augmentation
The quantity and quality of data play a crucial role in deep learning, especially in light of the empirical evidence supporting scaling laws. Unlike conventional cross-modal visual-language retrieval, where image-text pairs or video-text pairs can be readily collected from the web, CMR requires training data in the form of triplets, e.g., {reference image, relative text, target image}. In such triplets, the reference image and relative text jointly represent the user’s retrieval intent, with the relative text specifying flexible modification requests grounded in the rich visual objects and attributes of the reference image. The target image serves as the desired output of the retrieval process, with the relative text effectively describing the semantic differences between the reference and target images. However, in practice, triplets that capture such nuanced semantic differences are scarce, making constructing large-scale, high-quality triplet data a fundamental challenge in this field.
Early works primarily relied on manual annotation to construct triplet datasets. While these efforts provided a foundation for dataset construction, the manual annotation process is time-consuming and labor-intensive, leading to several significant limitations: (1) Limited Data Scale: The size of datasets is constrained by the human effort required for annotation; (2) Restricted Data Domains: Datasets are often confined to specific domains, limiting their generalizability; (3) Simplistic Relative Text Descriptions: The modifications expressed in the relative text are frequently simplistic, lacking the complexity required for real-world applications; (4) Annotation Noise and Uncertainty: Human annotations may introduce noise, which becomes a bottleneck for further advancements in the field.
To mitigate these issues, current work focuses on leveraging image and text generation techniques to automate the construction of triplet data. These approaches aim to improve the quantity of triplet data while having high-quality triplet annotations by minimizing the introduction of noise and uncertainty. Below, we highlight related key works and their technical innovations.
3.1.1 Automatic Data Construction
Works in this research line can be divided into two major categories according to the type of generation modality: (1) generating language content only  [1, 2, 3, 4, 5, 6, 7]; and (2) generating language and visual content simultaneously [8, 9, 10].
Generating Language Content: The augmentation of triplet data through text generation can be categorized into two main approaches: 1) generating text from a single image and 2) generating text describing differences between two images. These methods generate relative descriptions from a single image, typically by modifying its caption or masking parts of the input information. Liu et al. [1] develop a pipeline using large-scale image-text datasets (e.g., Laion-COCO) to generate relative descriptions by editing original captions. This is achieved using predefined templates (e.g., quantity changes, comparisons) or large language models (LLMs) like ChatGPT. The generated text is compared with candidate image descriptions to select the most similar target image, forming triplets. Two datasets are constructed: Laion-CIR-Template (template-based) and Laion-CIR-LLM (LLM-based), each containing approximately 16K triplets. The authors note that template-generated relative descriptions tend to be more effective, especially in fine-grained retrieval tasks, compared to those generated by LLMs. Hou et al. [2] propose a masked training strategy, which involves randomly masking patches of an image to create a reference image, while the original image serves as the target. An image caption generator is then used to produce a caption for the original image, which acts as the modification text, creating pseudo triplets. Zhang et al. [3] focus on generating challenging negative examples. They modified query sentences or replaced reference images with semantically mismatched alternatives to create pseudo triplets with subtle differences. This improved the model’s ability to distinguish examples and learn a better metric space for complex retrieval tasks. Liu et al. [4] introduce bidirectional training, where the model is trained on both forward (reference image + modification text to target image) and backward (target image + reverse modification text to reference image) queries. This expands data volume and strengthens multimodal fusion capabilities.
These methods generate descriptions of differences between two images to form triplets. Jiang et al. [5] propose a synthetic label generation strategy that consists of three steps: (1) selected visually similar image pairs, (2) generated difference descriptions for the pairs using pre-trained visual-language models (e.g., BLIP) and LLMs, and (3) filtered the generated text for semantic quality. This ensured high-quality triplet construction. Jang et al. [6] propose the Visual Delta Generator (VDG), which first searches for reference images and their corresponding target images in auxiliary datasets, ensuring that the image pairs have both similarities and differences. Then, the VDG generates text descriptions of the visual differences between the reference and target images, constructing pseudo triplets. Levy et al. [7] introduce LaSCo, a large-scale dataset that is ten times the size of existing CMR datasets. They selected visually similar image pairs with differing answers to the same question and used GPT-3 to generate transition text describing the changes. This approach resulted in a large-scale dataset LaSCo, significantly expanding available triplet data and boosting model performance.
Generating Language and Visual Content: Some works not only generate text but also generate images to further augment the triple data. Zhang et al. [8] propose CTI-IR, a GAN-based network that jointly trains generative and retrieval models to handle image generation and retrieval tasks. The model uses global-local collaborative discriminators to ensure semantic consistency between the generated image and the modification text by capturing both overall and fine-grained differences. Gu et al. [9] introduce a diffusion-based model for constructing a large-scale synthetic dataset called SynthTriplets18M. By leveraging existing image-caption datasets (e.g., COYO 700M, LAION-2B-en-aesthetic), they first generate diverse modified texts by replacing object terms in the original captions or editting operation via LLMs. This process creates a wide range of semantic variations, covering more visual concepts. Then, a latent diffusion model is used to generate high-quality target images based on the reference image and modification text. This approach enhances the model’s ability to handle diverse modifications and introduces capabilities like handling negative text and image mask conditions. Zhou et al. [10] propose one strategy, generating multi-modal triplets from large-scale image-text datasets. They use LLMs (GPT 3.5) to generate modification texts, which can be to change certain objects, colors, positions, etc, and then employ stable diffusion to create target images, ensuring alignment between the modified text and the generated image.
3.1.2 Noise/Uncertainty in Data
In addition to the volume of data, ensuring data quality is a critical focus in existing research. High-quality triplet data can significantly enhance model training effectiveness by mitigating the impact of noisy data, thereby improving retrieval performance. Data uncertainty and noise primarily manifest in two forms: (1) Content noise, where erroneous content is incorrectly deemed semantically relevant [5, 1, 9, 11, 2, 7]. For instance, the target image does not fully align with the retrieval intent specified by the modification text. This issue is particularly pronounced in automatically generated data. (2) Annotation uncertainty, where the target image in the triplet is not the sole positive example, as other images in the dataset may also satisfy the retrieval intent defined by the reference image and modification text [12, 13, 14, 15, 16, 17]. Existing research has addressed these challenges through a series of studies. Below, we review and summarize these efforts in detail.
Content Noise: Existing methods primarily employ threshold-based filtering strategies to ensure semantic relevance and consistency. These strategies focus on measuring the similarity between reference images, modified texts, and target images to align with the intended modification intention. Jiang et al. [5] and Liu et al. [1] both utilize semantic similarity filtering to ensure consistency between generated modification texts and image pairs. Jiang et al. [5] screen modification texts within the language space, retaining triplets with high semantic similarity to both reference and target images. Similarly, Liu et al. [1] generate modification instructions using a large language model, followed by sentence transformation to produce edited descriptions. They apply a similarity-based filter to select triplets with high semantic relevance between edited descriptions and reference images. Gu et al. [9] design a CLIP-based filtering strategy, calculating CLIP similarity between reference and generated target images, as well as between original and modified texts, and applying thresholds to ensure visual and semantic consistency. They introduce directional CLIP similarity to evaluate alignment between image and text modifications and verify that generated images reflect specified keywords or instructions.
Lin et al. [11] introduce a similarity-based data augmentation method. Specifically, they utilize CLIP’s image encoder to extract features from all images in the training set, and then compute the cosine similarity between the reference or target images and other images in the training set. Based on the similarity scores, several of the most similar images are selected as new reference or target images, thereby generating pseudo-triplets. To ensure quality, the generated pseudo-triplets select the most similar images, constrained by predefined similarity thresholds and a limit on the number of pseudo-triplets per image to control quality and prevent overfitting. Hou et al. [2] introduce a random sampling strategy based on the 3-$\sigma$ rule to select challenging samples for fine-tuning. They generate pseudo-modification texts for unlabeled image pairs, calculate distance metrics, and define a candidate range using the 3-$\sigma$ rule. Randomly selected samples from this range mitigate noise while ensuring sufficient challenge, enhancing model robustness. Levy et al. [7] develop an analytical tool to detect redundant noise in queries by evaluating the contribution of different modalities. Queries with minimal impact from either image or text are identified as containing redundant information. By identifying and filtering out triplets that include such redundant information, the quality of the generated data can be enhanced.
Annotation Uncertainty: To address the issue of annotation uncertainty in triplets, where a single query may correspond to multiple valid target images, existing methods focus on mitigating false positives and enhancing semantic diversity. These approaches can be categorized into two main strategies: uncertainty-aware modeling and constructing evaluation datasets with multiple ground-truth labels.
Some works focus on modeling uncertainty in data. Wen et al. [12] propose a target similarity distribution-based matching regularization module to handle false positives. This module calculates the similarity between the ground-truth target image and other candidates within a mini-batch, dynamically adjusting match scores to allow flexible handling of partially matching candidates rather than treating all non-target images as negatives. Zhang et al. [13] address triplet uncertainty by introducing the Consensus Network (Css-Net), which employs four distinct compositors to generate diverse image-text embeddings. This reduces reliance on any single potentially biased compositor and promotes consistent outputs through Kullback-Leibler divergence loss. In evaluation, decisions from the compositors are combined via a weighted scheme to enhance consistency. Chen et al. [14] propose a rank-aware uncertainty method to model many-to-many correspondences. Their approach includes three modules: (1) In-sample Uncertainty, which represents reference and target image features as Gaussian distributions to generate multiple potential matching candidates; (2) Cross-sample Uncertainty, which identifies potential positive samples by comparing similarity distributions across different samples; and (3) Distribution Regularization, which minimizes discrepancies between source and target feature distributions to improve cross-modal alignment. Chen et al. [15] further enhance uncertainty modeling by introducing fluctuations in the feature space that simulate multi-granularity queries. The model uses one-to-one matching, while for coarse-grained queries, it employs one-to-many matching within an uncertainty range. An uncertainty regularization module dynamically adjusts matching targets based on query precision, reducing false positives and improving retrieval accuracy.
Some works propose new evaluation datasets that annotate multiple ground truth labels for each query. Baldrati et al. [16] introduce the CIRCO dataset, which addresses the limitation of single ground truth labels by annotating multiple valid target images for each query. Using multimodal retrieval models, the top 100 relevant images are retrieved for each query, filtered to 50 visually similar images, and annotated to include all valid matches. This results in an average of 4.53 ground truth labels per query, reducing false negatives and better simulating real-world complexity. Similarly, the CFQ (Challenging Fashion Queries) dataset [17] introduces multi-label annotations, where each query includes multiple positive and negative labels. Annotators select target images from the same category that match the query as positive labels and those that do not as negative labels. This provides a more comprehensive and challenging benchmark for evaluating text-guided fashion image retrieval models.
3.2 Model Architecture
For input reference images and modification texts, the core challenge of combined multimodal retrieval lies in how to integrate the two modalities, that is, to represent a composite semantics that is similar to the reference image but adjusted in detail according to the modification text. Existing works can be categorized into three main aspects: (1) The most common approach is to design feature combiners for visual and language modalities, fusing the two modalities to achieve modified multimodal semantic representation; (2) Plug-and-play components, which offer broad compatibility across different existing methods and hold significant summarizing value; (3) Non-combiner methods that take an alternative path, implementing combined multimodal retrieval through novel architectures, providing important inspiration.
3.2.1 Design of Visual-Linguistic Combiner
Based on the key highlights of different combiner methods, ranging from simple to complex design structures, we classify them into the following 6 categories: (1) Global-level combination based on coarse-grained modality features; (2) Local-level combination focusing on fine-grained modality features; (3) Hierarchical combination that incorporates multiple feature granularities; (4) Combination that models structured relationship information between modality samples; (5) Combination that leverages external knowledge for enhancement; (6) Combination that considers data polysemy by adopting ambiguity probability modeling. Below, we introduce the details of each category.
Global-level Multimodal Combiner: Various methods have been proposed to effectively compose the global-level visual reference and textual modification into new semantic representations that resemble the visual target. The popular approach is based on gating and residual connection mechanisms. Vo et al. [18] introduce the Text Image Residual Gating (TIRG) method, which employs a gating mechanism and residual connection to modify image features based on text input while retaining the original image features when text modifications are insignificant. The gating mechanism computes a weighted sum of the gate and residual features, where the gate feature preserves the query image, and the residual feature introduces text-specific modifications. Dodds et al. [17] introduce a Residual Attention Fusion (RAF) mechanism for text-guided fashion image retrieval. RAF calculates attention weights from text features, applies them element-wise to image features, and uses residual connections to maintain alignment in the embedding space. This approach achieves global modifications while preserving the primary structure of the image features. Baldrati et al. [19] design a Combiner network that maps image and text features into a shared embedding space using a pre-trained CLIP model. The Combiner processes concatenated features through linear layers, ReLU activations, and Dropout to generate a combined feature. In the subsequent works [20, 21], they fine-tune CLIP’s text encoder for task-specific adaptation and propose a Combiner that outputs a normalized combination of text and image features, including convex and learned mixtures. Lin et al. [11] also fine-tune CLIP’s encoders and employ a fusion module with weighted summation, concatenation, and bilinear pooling to capture complex modification intents.
Besides, Lee et al. [22] present the Content-Style Modulation (CoSMo) framework, which decomposes textual features into content and style components. Content modulates spatial activation patterns of image features, while style adjusts their distribution via learnable affine transformations. An adaptive fusion module dynamically balances these components based on the text, optimizing multimodal feature integration. Anwaar et al. [23] design a method based on complex space mapping and symmetry constraints. Specifically, they map image and text features into a complex space. Text features define rotational transformations from source to target image features, with a rotational symmetry constraint, ensuring that multiplying the target image feature with the complex conjugate of the query text feature yields a representation similar to the query image feature. Deep metric learning is used to optimize the composition and mapping of these complex vectors. Xu et al. [24] introduce a composition-and-decomposition framework. The composition network couples reference image and text features into a joint representation, while the decomposition network decouples the target image into visual and textual subspaces. An attention mechanism selects key components, and techniques like bilinear pooling enhance feature extraction, forming a closed-loop system for mutual optimization.
Local-level Multimodal Combiner: Current local feature-based CMR methods can be primarily categorized into three approaches: attention-based feature interaction methods [25, 26, 27] that achieve fine-grained correspondence through multi-layer or bidirectional attention mechanisms; semantic enhancement and suppression methods [28, 29, 30] that focus on selective modulation of specific semantics via dynamic weights or kernels; and concept alignment and distribution learning methods [31, 32] that emphasize visual-semantic alignment at the conceptual level and optimize matching boundaries. While these approaches differ in their technical strategies, they all aim to improve the utilization of local features for more precise retrieval performance.
Attention-based feature interaction methods focus on establishing fine-grained correspondences between image regions and text words. For instance, Hosseinzadeh and Wang [25] employ a multi-layer bidirectional attention mechanism to enhance the understanding of specific image region modifications in relation to text instructions. Each layer’s two sub-modules compute linguistically attended visual features and visually attended linguistic features, creating a detailed interaction between the visual and textual features and gradually generating richer joint representations. Similarly, FashionViL by Han et al. [26] introduces a dynamic, learnable attention module to build region-word correspondences, further refining feature fusion through bilinear pooling, which captures second-order statistics to better integrate complex visual and textual features. Zhang et al. [27] also propose to achieves deep integration of local modal features through Multi-modal Complementary Fusion (MCF) and Cross-modal Guided Pooling (CGP) modules. The MCF module uses a double-interaction mechanism to compute attention weights and interaction maps between the reference image and text, projecting text features into the visual space and fusing them with image features via a modified gating operation to enrich the image representation with key textual information. The CGP module explores both intra-modality and inter-modality relationships, calculating cross-modal relevance between local image features and text word vectors, as well as intra-modal interactions among channels of the reference image. It dynamically assigns weights to local features, generating a holistic representation that integrates cross-modal information. This design emphasizes leveraging local modal features for finer-grained interactions and more accurate retrieval outcomes.
Semantic enhancement and suppression methods, on the other hand, center around selectively enhancing or suppressing specific semantic features of the image based on textual information. Pang et al.’s Weak-text Semantic Augment (WSA) module [28] enhances short or incomplete text queries by incorporating relevant image information. WSA expands the text feature dimensions and uses a multi-layer perceptron to map image features into the text’s latent space. An attention mechanism then integrates these image features and maps them into the semantic space to enhance the text’s semantic expression. This allows WSA to flexibly select high-level image features pertinent to the text, improving the understanding of user query intent, especially for brief or incomplete text descriptions, thus enhancing retrieval accuracy and robustness. Zhang et al.’s approach [29], involving multi-semantic dynamic suppression and text-semantic complementary selection, divides the retrieval task into two sub-processes: suppressing irrelevant image details and enhancing relevant ones based on the text, ensuring that the final image representation aligns closely with the textual description. Similarly, the SAC framework [30] proposed by Jandial et al. uses a two-stage approach, first identifying salient regions in the image relevant to the text and then adjusting these regions while preserving other areas of the image, employing a combination of LSTM and batch normalization for feature fusion. These methods are united by their focus on the selective enhancement or suppression of semantic features, using dynamic weights or kernels to modulate feature importance while maintaining the integrity of unmodified areas.
Finally, concept alignment and distribution learning methods emphasize aligning visual and textual concepts at the semantic level. Zhao et al. [31] achieve concept-level visual-semantic alignment. By concatenating feature tokens of the reference image and target image, and then applying a Transformer to perform context modeling on the concatenated features, the model subsequently employs an attention-based multiple-instance learning framework to compute alignment scores between visual concept features and semantic concept embeddings in the text modifiers. This alignment is then used to guide the fusion of the reference image and text features, improving retrieval accuracy. Similarly, Wang et al. [32] explore the positive and negative correspondences of local features, introducing a threshold optimization mechanism to dynamically distinguish between matched and unmatched samples. By optimizing this threshold, the module strengthens the model’s ability to distinguish between local relevances and irrelevances, enabling it to better handle highly similar samples and avoid mismatches.
Hierarchical Multimodal Combiner: The design of hierarchical multi-modal combiner is driven by the need to capture multi-granular visual and textual relationships, enabling a more comprehensive fusion of information across different levels of abstraction. Existing methods can be categorized into two primary types based on how hierarchical features are generated: encoder multi-layer strategies [33, 34, 35, 36, 37, 38, 39, 40] and global-local feature strategies [41, 42, 43, 44, 29].
The first type approaches utilize features extracted from different layers of encoders, such as CNNs or Transformers, allowing for a more detailed representation of image-text interactions. Chen et al. [33] propose a method where visual features from multiple layers of a CNN are concatenated with textual features, forming a visiolinguistic representation. This representation is then processed through two attention mechanisms: a self-attention stream that uncovers region-to-region dependencies and a joint-attention stream that modulates the preservation of visual content based on the textual input. The use of features from different CNN layers enables the model to preserve both high-level semantics and low-level visual details, enhancing the overall multigranular fusion of information. Dodds et al. [34] propose a multimodal retrieval model that extracts image features from different ResNet50 layers, treating them as tokens representing various parts of the image. These tokens, along with tokens generated from textual features, are concatenated into a unified sequence. Through self-attention, the model enables interaction between image and text tokens, facilitating cross-modal fusion. Using features from different ResNet layers, the model captures both global and local visual characteristics, making it effective for retrieving target concepts from modified query images.
Similarly, Jandial et al. [35] present a framework that extracts multi-scale visual features from a CNN, forming a visual feature pyramid. Each level of this pyramid is aligned with textual embeddings generated by a GRU, and these multimodal embeddings are further processed through a Semantic Feature Transformation module. This module uses attention mechanisms to align the visual and textual information at each level, followed by a Hierarchical Feature Aggregation module that aggregates these multi-level representations. This hierarchical processing ensures a comprehensive visiolinguistic composition, allowing the model to address the complexities of text-conditioned image retrieval. Yang et al. [36] introduce a model that integrates modification and auxiliary relation modeling, using a multi-level aggregation method to fuse visual and textual information. Through the Hierarchical Aggregation Transformer, the model captures both low-level texture details and high-level semantic information, ensuring a richer representation for multi-modal composition. This design enhances the model’s flexibility in adapting to different types of query texts and improves the retrieval process by ensuring a more accurate alignment between reference images and target images.
These methods leverage hierarchical feature extraction to further localize the modified specific image regions. Zhang et al. [37] focus on image localization for composed image retrieval, employing a multi-level approach that leverages feature maps from shallow, middle, and deep layers of the CNN. The localization process is applied across these layers, using up-sampling and down-sampling operations to align the localization distribution, resulting in an improved fusion of multi-modal features. This approach enhances the accuracy of localizing relevant areas of the image in response to textual input, thereby supporting precise multi-modal feature integration. A similar multi-level design is evident in Xu et al.’s approach [38], where hierarchical feature maps from a Vision Transformer are used to capture visual details at various levels. Then, the designed global-local alignment bridges the gap between composed queries and target images by focusing on both global embedding spaces and local region details through spatial attention mechanisms. This design improves the model’s ability to refine specific areas of the reference image according to the modification text, enhancing retrieval accuracy. Li et al. [39] extend this idea with a method that extracts multi-grained visual and textual features from different layers of a pre-trained ResNet50. By incorporating a Multi-Grained Semantic Construction module, they capture fine, mid, and coarse-grained information from both modalities. Their attention mechanism, supervised by a triplet ranking loss, ensures that textual features focus on the modified regions while excluding the preserved ones, optimizing the multi-modal fusion process. Besides, Huang et al. [40] introduce the language-guided local infiltration system, which focuses on hierarchical feature extraction and fusion to modify specific image regions while preserving the overall structure. Specifically, The language prompt visual localization module generates a localization mask to identify the target region for modification based on the text, enabling precise control over which part of the image is altered. The text infiltration with local-awareness module then applies the mask to adjust the image, ensuring that local changes align with the text while maintaining the integrity of other regions.
The second type approaches emphasize a hierarchical approach to feature extraction, where local, regional, and global information is progressively integrated, ensuring a rich representation for multimodal retrieval tasks. Wen et al. [41] propose a method that incorporates both local and global composition modules to handle multimodal queries. The local composition module focuses on fine-grained visual descriptors, while the global composition module integrates global visual features with textual modifications. These modules use attention mechanisms to align the image and text, allowing the model to capture detailed local interactions as well as broader semantic shifts. A mutual enhancement module is introduced to facilitate knowledge sharing between the two composition modules, reinforcing the performance of both. This hierarchical design enables the model to handle a wide range of modification types, from specific local changes to more abstract visual adjustments, by combining these different levels of information. Liu et al. [42] introduce a model that leverages a pre-trained vision-and-language (V&L) model, OSCAR, to map composed queries consisting of a reference image and textual description into an embedding space. While OSCAR processes local text sequences, the model adapts to images by using a pre-trained ResNet to extract global image features. This design allows for the combination of global image features with modification text, processed through a multi-layer transformer structure with self-attention mechanisms. By using global image features instead of region-specific representations, the model captures a broader understanding of the image and text relationship, while still leveraging local text sequence processing.
Goenka et al. [43] introduce a fashion image retrieval model that extracts features at multiple levels: the global image, cropped clothing regions, regions around fashion landmarks, and areas identified by a pre-trained object detector (RoIs). These features are then integrated with textual feedback through a Vision-Language Transformer, while the target image processing focuses solely on visual features to reduce inference costs. This method enhances the fusion of visual cues and textual information, facilitating efficient retrieval. Wen et al. [44] design a multi-granularity feature processing mechanism using fine-grained local and global composition modules. Their model assigns specific matching tokens for attributes like color, belt, and collar, which adaptively aggregate features during training. This approach enables the effective use of multimodal data, from fine-grained details to coarser, more abstract information. Zhang et al. [29] integrate CNNs with Transformers for image feature extraction, where the CNN captures local features while the Transformer models global edge information. This hybrid approach improves the representation of reference images by combining local and global features, while LSTMs handle textual feature extraction. The final fusion of these features through residual and gating mechanisms enhances the model’s ability to match query-image and target-image pairs.
In summary, these methods emphasize hierarchical feature extraction from various encoder layers to effectively capture both low-level and high-level visual and textual information. By incorporating attention mechanisms and leveraging different feature granularities, these methods improve the flexibility and precision of multi-modal composition, leading to more accurate retrieval and understanding of composed queries.
(4) Structured relational multimodal combiner: Some studies collectively advance the field of multi-modal composed retrieval, from foundational hierarchical composition learning to innovative applications of residual learning, comprehensive relationship reasoning, complementary association learning, and style-guided patch interaction. While these research methods focus on different technical strategies, they all aim to enhance the compositional modeling of multi-modal information from various perspectives, improving the retrieval performance of complex composed queries.
Xu et al. [38] introduced a Hierarchical Composition Learning (HCL) framework to tackle hierarchical composition problems. This method progressively integrates complex relationships between text and images through a multi-level compositional structure. Similarly, Li et al. proposed the Residual Learning for Text and Image Composition using Graph Convolutional Network (RTIC) method [45], introducing residual connections to enhance information flow between text and images and improve retrieval accuracy and robustness. RTIC incorporates a residual learning module based on GCNs, which accumulates and propagates differences between modalities through multiple residual connections, and an adaptive weighting mechanism to dynamically adjust the importance of different layers.
Zhang et al. [46] developed comprehensive relationship reasoning, which utilizes explicit graph structures to capture intricate relations among elements in text and images. The approach dynamically constructs and updates graph structures to accommodate various querying needs. Moreover, Jiang et al. [47] presented Complementary Association Learning (CaLa), emphasizing the importance of mining complementary information from texts and images. CaLa employs a dual-stream encoder with cross-attention mechanisms to enhance retrieval performance. Most recently, Chen et al. introduced the Style-guided Patch Interaction (SPIRIT) method [48], addressing text-feedback-based image retrieval in the fashion domain by enhancing interactions between image patches through the incorporation of style information. SPIRIT utilizes a style encoder to extract style features from input images, employs a style-guided attention mechanism to adjust the interaction strength between image patches dynamically, and incorporates a text feedback module to further optimize retrieval results based on user-provided text input.
(5) Knowledge-enhanced multimodal combiner: Several approaches have been proposed to enhance the performance of multi-modal compositional retrieval by leveraging external knowledge or specific mechanisms. Among these, Zhang et al. [49] emphasizes the modeling of geometric structures and cross-modal relationships. Moreover, Zhang et al. [3] and Liu et al. [50] further improve upon this by introducing external knowledge and sophisticated negative sample generation strategies to optimize multi-modal fusion. In contrast, Bai et al. [51] focuses on utilizing sentence-level prompts to improve retrieval accuracy and generalization capability. Specifically, Zhang et al. [49] proposed a Geometry-Sensitive Cross-Modal Reasoning network (GSCR), which achieves joint modeling of the visual-semantic relationships and geometric structures of images with modified texts through a Geometry-Sensitive Inter-Modal Attention module (GS-IMA) and a Text-Guided Visual Reasoning module (TG-VR). While GS-IMA explicitly and implicitly incorporates spatial structure into cross-modal attention, TG-VR is utilized to model semantic information not present in the reference image, guiding further visual reasoning. This design emphasizes the importance of leveraging the spatial structures of images and enhancing cross-modal interactions when dealing with complex queries. Similarly, Zhang et al. [3] introduced Multi-modal Compositional Example Mining (MCEM), an approach that optimizes fusion across different modalities by generating component-level negative samples. By pairing logically mismatched sentences with reference images and introducing a novel sentence augmentation method to generate more challenging multi-modal negative samples, MCEM assists the model in learning a better metric space. This method highlights the importance of creating sophisticated negative samples to improve the understanding of multi-modal data by the model, showcasing innovation, particularly in utilizing external knowledge. Building upon this foundation, Liu et al. [50] presented a Dual Multi-modal Encoder-based Re-ranking method (DMER), which uses two independent multi-modal encoders to encode images and text separately, then re-ranks candidate images based on their similarity scores. By integrating an external knowledge graph, DMER enables the model to leverage additional semantic information for reasoning, thereby enhancing its expressiveness. This approach underscores the value of utilizing an external knowledge graph to supplement traditional single-encoder methods, demonstrating significant effectiveness especially on large-scale datasets. Bai et al. [51] introduced Sentence-Level Prompts Benefiting Composed Image Retrieval (SLPCIR), aiming to enhance CIR performance through sentence-level prompts. By parsing textual descriptions to identify key sentence constituents and generating corresponding prompts for each constituent, SLPCIR guides the model in more accurately locating and retrieving target images. This method illustrates how sentence-level prompts can be used to improve retrieval accuracy, showing strong generalization capabilities particularly in zero-shot and few-shot scenarios, thus enhancing the effectiveness of multi-modal compositional retrieval.
(6) Ambiguous uncertainty/probabilistic multimodal combiner: Neculai et al. [52] addressed the challenge of combining multiple query inputs in multimodal image retrieval by proposing a Multimodal Probabilistic Combiner (MPC). Existing approaches are typically limited to handling single or dual query inputs, which constrains their capability to cope with complex scenarios. The MPC is designed to retrieve target images based on an arbitrary number and types of queries. To this end, the authors established a new benchmark using the MS-COCO dataset to evaluate the performance of multimodal image retrieval under varying modalities and query counts. The MPC learns probabilistic embeddings and employs parameterized probabilistic combination rules to synthesize these embeddings, resulting in composite multivariate Gaussian distributions that capture both the semantics and uncertainty of the input data. Xu et al. [53] focused on addressing the issues of semantic polysemy and data uncertainty. Traditional one-to-one mapping methods struggle to effectively handle these challenges due to their inability to adequately capture the diversity of multimodal data. To tackle this issue, the authors introduced the Set of Diverse Queries with Uncertainty Regularization (SDQUR). SDQUR leverages the Q-Former from BLIP2 to extract multiple deterministic embeddings from unimodal or multimodal inputs, where each embedding represents a potential matching factor. To address data uncertainty caused by label noise, SDQUR incorporates an uncertainty regularization module that encodes the input as a distribution, allowing for the sampling of multiple instances to establish probabilistic many-to-many correspondences.
3.2.2 Design of Plug and Play Structure
Based on the key highlights of different approaches in achieving plug-and-play structures, we classify them into the following four categories: (1) Modular design to enhance plug-and-play capability [54, 55]; (2) Dynamic configuration and adaptive mechanisms [56, 57]; (3) Data alignment and feature fusion techniques [28, 46]; (4) Training strategies and re-ranking improvements. Below, we introduce the details of each category [50, 4].
(1) Modular design to enhance plug-and-play capability: To improve the flexibility and scalability of multi-modal retrieval systems, some methods adopt modular network architecture designs, allowing these solutions to be easily integrated into existing systems. Chen et al. [54] proposed a Joint Visual Semantic Matching (JVSM) model that associates visual and textual modalities by constructing a shared discriminative embedding space, enabling image-text matching. The JVSM employs a four-module architecture (visual embedding, text embedding, semantic projection, and combination module), optimized with various loss functions, allowing it to be seamlessly integrated into existing systems without major architectural changes. Experimental results demonstrate its superior performance in handling complex queries, showcasing strong plug-and-play capabilities. Yang et al. [55] focused on addressing asymmetries and heterogeneities between cross-modal data, proposing a novel alignment loss function to support effective feature mapping and alignment. This method is designed as an independent component that can be directly applied to existing CMR systems, enhancing the association between different modalities and improving system flexibility and adaptability.
(2) Dynamic configuration and adaptive mechanisms: Some works emphasize dynamic adjustment capabilities based on specific application scenarios, allowing systems to flexibly respond to different query types and user needs. Kim et al. [56] developed a dual compositional learning framework capable of dynamically adjusting composition and fusion strategies according to specific application scenarios, adapting to changes in user queries, thus significantly enhancing interactive image retrieval performance. This flexible design enables systems to quickly adapt to new environments or requirements, demonstrating high flexibility and ease of integration. Zhu et al. [57] introduced an adaptive multi-expert collaborative network that selects the most suitable expert models for processing inputs, providing more personalized and precise retrieval services. This dynamic selection mechanism based on input characteristics allows to rapidly adapt and integrate into various text-guided image retrieval systems, greatly enhancing system functionality and robustness.
(3) Data alignment and feature fusion techniques: Addressing asymmetries and heterogeneities between heterogeneous data, researchers have proposed various feature alignment and fusion techniques to enhance cross-modal retrieval performance. Pang et al. [28] presented a framework that combines multiple types of features and ensures accurate correspondence through cross-modal alignment algorithms. The authors designed a multi-layer perceptron for fusing heterogeneous features and used contrastive loss for aligning different modalities. This method not only improves the relevance and accuracy of retrieval results but also demonstrates its potential as a plug-and-play component. Zhang et al. [46] proposed a comprehensive relationship reasoning mechanism that analyzes relationships between images and texts in detail, constructing a refined relational graph and using graph convolutional networks to capture these relationships. This comprehensive reasoning approach can be directly applied to existing CMR systems, significantly enhancing the performance of composed query-based image retrieval.
(4) Training strategies and re-ranking improvements: Finally, several studies explored how to directly enhance retrieval system performance through improved training strategies and candidate set re-ranking. Liu et al. [50] utilized dual multi-modal encoders for the candidate set re-ranking, offering a convenient way to optimize final retrieval results. This lightweight design allows for rapid implementation without altering the original system architecture, thereby improving system efficiency and accuracy. Liu et al. [4] investigated a bi-directional training strategy via text prompt learning, emphasizing the importance of lightweight design, enabling quick deployment without changing the original system architecture, and further enhancing system understanding and retrieval capabilities. This highlights the advantages of bi-directional training in improving retrieval effectiveness and demonstrates its efficacy as a plug-and-play component.
3.2.3 Design of Non-combiner Architecture
Non-combinatorial structural design refers to methods that avoid directly merging or combining different modalities of data (such as text and images) when processing multimodal data. Instead, these approaches achieve cross-modal information fusion and utilization through indirect means, aiming to circumvent the complexity and limitations that traditional combinatorial methods may introduce. Specifically, non-combinatorial structural design attempts to preserve the uniqueness of each modality’s data through various strategies, and on this basis, conduct information interaction, enhancement, and complementation to achieve more effective information retrieval and utilization.
For instance, Zhang et al.[8] proposed the Tell-Imagine-Search framework, which consists of three modules. The Tell module generates a detailed text description based on the input text and reference image; the Imagine module synthesizes an imaginary image according to this description; and the Search module uses the synthesized image to search for the most similar real image in the database. This indirect cross-modal information fusion approach improves retrieval accuracy and enhances the system’s creativity and flexibility. Subsequently, Delmas et al.[58] proposed the ARTEMIS framework, which utilizes attention mechanisms to enhance cross-modal retrieval performance. ARTEMIS allows the model to simultaneously consider explicit matching of text content and implicit similarities between image features, introduces a new loss function to guide learning, and adopts a lightweight network architecture to ensure inference efficiency, reflecting the preservation and utilization of each modality’s uniqueness.
Fu et al.[59] constructed a Multi-Order Adversarial Representation Learning (MOARL) method to address representation learning. MOARL includes pixel-level, feature-level, and semantic-level multi-order adversarial training, balancing the learning process at different levels through specially designed loss functions, showcasing its non-combinatorial structural characteristics and improving the model’s understanding of complex queries. Han et al.[60] proposed the FAME-ViL multi-task model, which employs a shared encoder-decoder structure to tackle the challenges of heterogeneous vision-language tasks in the fashion domain. The encoder extracts universal vision-language features, while different decoders correspond to specific tasks such as product classification, attribute prediction, and image retrieval. This architecture, which processes multimodal information in stages, demonstrates the flexibility and efficiency of non-combinatorial structural design. Yang et al.[61] proposed the Decompose Semantic Shifts (FSS) method to address the semantic shift problem. FSS identifies key query elements that may cause semantic shifts through semantic parsing and learns stable representations of these elements through a contrastive learning framework for precise image retrieval. This showcases the potential of non-combinatorial structural design in enhancing the retrieval relevance and robustness of complex queries. Wen et al.[62] transferd multimodal fusion to the raw data level, leveraging vision-language pre-training models to achieve multimodal encoding and cross-modal alignment. This framework resolves the embedding space shift problem that may arise from feature-level fusion, further validating the effectiveness of non-combinatorial structural design.
3.3 Loss Optimization
For the combined features of reference images and modified text, achieving constraints between these and the target images is a critical issue in the field. The loss optimization designs in existing methods mainly focus on three aspects: (1) fundamental contrastive learning constraints; (2) additional consistency constraints to refine details; and (3) joint optimization constraints that integrate multi-task learning.
3.3.1 Based on Contrastive Learning
The two most common contrastive learning losses are batch-based classification loss and soft triplet-based loss, both variants of Softmax Cross-Entropy Loss [18]. First, let us define some key variables. Assume we have a mini-batch of training triplet data containing $B$ queries, where each query composed of the reference image and modification text is denoted as $\psi_{i}$ and the target image is denoted as $\phi^{+}_{i}$. The negative samples can be randomly selected by $K$-1 target images of other queries from the same mini-batch, denoted as $(\phi^{-}_{1},\phi^{-}_{2},\ldots,\phi^{-}_{K-1})$. These negative and positive samples form a set $N_{i}$ for computing the loss function. Below, we introduce these two losses and their relationships.
Softmax Cross-Entropy Loss [18]: The goal is to pull the ”modified” query features closer to the target image features while pushing away dissimilar image features. Specifically, the loss function is formulated as:
$L=\frac{1}{MB}\sum_{i=1}^{B}\sum_{j=1}^{M}-\log\frac{\exp\{\kappa(\psi_{i},\phi^{+}_{i})\}}{{\sum_{\phi_{i,j}\in N_{i}}\exp\{\kappa(\psi_{i},\phi_{i,j})\}}},$ (4)
where $\kappa$ is a similarity kernel function (e.g., the dot product or negative L2 distance). The maximum value of $M$ is $\binom{B}{K}$, but it is often selected as a smaller value for tractability. This loss ensures similar samples are close in the embedding space, while dissimilar ones are pushed apart.
Batch-based Classification Loss [63, 64, 61, 57]: When the number of negative samples $K$ equals the batch size $B$, the above loss simplifies to:
$L=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp\{\kappa(\psi_{i},\phi^{+}_{i})\}}{{\sum_{j=1}^{B}\exp\{\kappa(\psi_{i},\phi_{i,j})\}}}.$ (5)
In this setting, each query is compared against all other samples in the mini-batch, making the loss function more discriminative and accelerating convergence. However, it also increases the risk of overfitting.
Soft Triplet-based Loss [42, 25, 55]: In the case of using the minimum value $K=2$, each query has only one negative sample, the loss simplifies to:
$L=\frac{1}{MB}\sum_{i=1}^{B}\sum_{j=1}^{M}\log{(1+\exp{(\kappa(\psi_{i},\phi^{-}_{i,j})-\kappa(\psi_{i},\phi^{+}_{i}))})}.$ (6)
This form relaxes the strict distance constraints between positive and negative samples, making the model more stable and flexible for complex data distributions. As demonstrated in [18, 25], selecting an appropriate loss function allows the model to achieve better performance across different datasets.
Triplet Ranking Loss [54, 33, 49]: Unlike soft triplet loss, which relaxes the distance constraint with a logarithmic function, triplet ranking loss uses a hard margin constraint. It constructs a triplet of an anchor (e.g., a multimodal composed query), a positive(e.g., a target image), and a negative sample (e.g., an unrelated image) to optimize the embedding space, ensuring similar samples are close and dissimilar ones are pushed apart. The triplet loss function is formulated as:
$L=[\kappa(\psi_{i},\phi^{+}_{i})-\kappa(\psi_{i},\phi^{-}_{i,j})+\gamma]_{+}$ (7)
where $[\cdot]_{+}=\max(0,\cdot)$, and $\gamma$ is a predefined margin. The loss occurs only when the distance to the positive sample exceeds that to the negative sample plus the margin. Soft triplet loss is more flexible and stable in complex data, while triplet ranking loss enforces stricter separation, making it suitable for clear sample distinctions. Though it converges faster, it may become unstable with complex or noisy data.
Bi-directional Triplet Ranking Loss [54]: Different from the conventional triplet loss which only constrains the positive combination $\psi^{+}_{i}$, it introduces the negative combination $\psi^{-}_{i}$, i.e., the reference image and modified text that are irrelevant to the target image, which can be expressed as:
$L=$ $[\kappa(\psi^{+}_{i},\phi^{+}_{i})-\kappa(\psi^{-}_{i},\phi^{+}_{i})+\gamma]_{+}+$ (8)
$+[\kappa(\psi^{+}_{i},\phi^{+}_{i})-\kappa(\psi^{+}_{i},\phi^{-}_{i,j})+\gamma]_{+}.$
Additionally, they also introduce a text-matching loss to optimize the alignment between the positive text $t^{+}_{i}$, i.e., the caption of the target image, and the negative text $t^{-}_{i,j}$, i.e., the caption of the irrelevant target image. Specifically, the loss function is formulated as:
$L=$ $[\kappa(\psi^{+}_{i},t^{+}_{i})-\kappa(\psi^{-}_{i},t^{+}_{i})+\gamma]_{+}$ (9)
$+[\kappa(\psi^{+}_{i},t^{+}_{i})-\kappa(\psi^{+}_{i},t^{-}_{i,j})+\gamma]_{+}.$
In addition to designing the loss function, existing work also focuses on the negative sample mining strategies and the similarity metric function.
Negative Sample Mining: As a fundamental role in training, there are two common strategies: mini-batch hardest negative mining [49, 65, 66] and mini-batch semi-hard negative mining [54, 33]. The former only selects negative examples with the greatest similarity to the combined query features, while the latter selects negative examples with a higher similarity to the combined query than the target image. In Work A, it is validated that semi-hard mining is more stable and faster converging than hardest mining.
Unlike simple selection, Zhang et al. [3] propose a method to enhance multimodal fusion by replacing parts of the query’s text to generate harder negative samples, improving model learning. They introduce two techniques for mining hard examples: (1) synthetic composition examples, which create negative samples by replacing the text while keeping the image part unchanged, forming a new sample as $f=r+t^{\prime}$, where $t^{\prime}$ is from another query. (2) augmented synthetic composition examples, which replace only parts of the text, generating even harder-to-distinguish negatives. A mask vector $s$ randomly determines which dimensions of the text to retain or replace, with a Bernoulli distribution governing the decision. The new text embedding is $\tilde{t}^{\prime}=s\odot t+(1-s)\odot t^{\prime}$. They also introduce a dynamic replacement ratio, sampling $p$ from a Beta distribution for more diverse hard examples.
To optimize the use of these hard examples, they propose an improved loss function:
$L=\frac{1}{B}\sum_{i=1}^{B}-\log\frac{\exp(\kappa(\psi_{i},\phi^{+}_{i}))}{\exp(\kappa(\psi_{i},\phi^{+}_{i}))+\sum_{j=1}^{B-1}\exp(\kappa(\phi^{+}_{i},f_{j}))}$ (10)
This formula ensures that the query $\psi_{i}$ has high similarity with the target image $\phi^{+}_{i}$ and low similarity with other synthetic composed negative samples $f_{j}$, enhancing the model’s discriminative power.
Similarity Metric Function: Conventional contrastive learning losses typically employ Euclidean distance or cosine similarity as metrics for measuring the distance between samples [18, 12, 10]. However, these methods consider only the absolute distance or angular relationship between samples, neglecting the combined effect of both factors. To address this limitation, Zhang et al. [29] propose the Triangle Area (TA) as a novel metric for sample distance measurement. Specifically, as illustrated in Figure A, for an anchor sample $a$, a positive sample $c$, and a negative sample $b$, the distances between the samples are represented by the areas of triangles $Oab$ and $Oac$, where $O$ denotes the origin. TA is defined as:
$\text{TA}(\psi_{i},\varphi_{i})=\frac{1}{2}|\psi_{i}||\phi_{i}|\sqrt{1-\left(\frac{\psi_{i}\cdot\phi_{i}}{|\psi_{i}||\phi_{i}|}\right)^{2}}$ (11)
TA offers two main advantages: (1) It considers both the absolute distance and angular relationship between samples, allowing the model to optimize both aspects without needing manual weight adjustments. (2) For smaller datasets, TA uses a squared distance, amplifying differences and slowing convergence to prevent overfitting. For larger datasets, it uses the area as a distance metric, avoiding excessive magnification of differences and enabling faster convergence and reduced training time.
3.3.2 Consistency Constraint
To ensure that visual and linguistic information can be effectively represented and mutually complementary within a shared space, researchers have proposed methods based on consistency constraints. These methods aim to capture the complementary information between images and text, enabling them to jointly express query intentions and accurately reflect user modification intentions, thereby generating more relevant retrieval results.
Works [35, 30, 23] focus on consistency loss for image and text reconstruction, aiming to construct a robust and meaningful joint representation space. Within this space, the model attempts to reconstruct the original input images and texts as faithfully as possible, learning their associations and semantic relationships to improve retrieval accuracy. Additionally, works [41] and [36] delve into the reconstruction consistency between global and local composite features. For instance, Wen et al.[41] utilize fine-grained local and global combination modules to achieve multimodal integration and introduce an inter-enhancement module to promote mutual improvement between local and global features. It enhances knowledge transfer across different feature levels using KL divergence loss and feature-level $L_{2}$-norm loss to optimize the consistency between local and global features. This approach not only improves the expressiveness of local features but also reinforces the synergy between the global and local combination modules.
Pang et al. and Zhang et al. [28, 27] introduce Relative Caption-aware Consistency (RCC) loss to address the heterogeneous semantic gap between images and texts. RCC promotes effective alignment and cross-modal fusion by minimizing the discrepancy between generated text snippets and original text snippets. Specifically, RCC employs an improved LSTM structure to generate, word by word, a textual description of the differences between reference and target images. It uses cross-entropy loss to measure the gap between the probability distribution of the generated text and the true labels. By minimizing this loss, the model adjusts its parameters during training to make the generated text closer to the user-provided modified text, thus achieving effective alignment of different modal features in a unified embedding space.
From the perspective of consistency constraints between multi-branch networks, works [13, 12, 61, 57] propose various mechanisms to ensure that different branches or expert networks produce consistent feature representations when processing the same input, thereby enhancing the model’s robustness and generalization capability. Zhang et al. [13] introduce the consensus network, which uses multiple diverse compositors to generate image-text embeddings, each providing a unique perspective. They incorporate KL consistency constraints through Kullback-Leibler divergence loss to promote mutual learning among different compositors, reducing individual biases and reaching consensus, which increases robustness against noisy annotations. Wen et al. [12] design a teacher-student framework with two branches: the teacher branch guides directly based on target-query relationships, while the student branch mimics the behavior of the teacher. They use KL consistency constraints to ensure that the student branch learns consistent conflict modeling from the teacher branch, reducing bias and improving the effectiveness of multimodal query combinations and candidate image ranking quality.
Yang et al. [61] propose a Semantic Shift Network (SSN), which decomposes text instructions into degradation and upgradation steps, explicitly handling the transition from reference images to target images. Through KL consistency constraints, SSN ensures that the positive sample feature distribution aligns with the target distribution while increasing the difference between negative samples and the target image, enhancing the accuracy of positive sample features and the model’s ability to distinguish between positive and negative samples. Zhu et al. [57] present an adaptive multi-expert collaborative network, which introduces multiple expert networks focusing on specific types of feature extraction, such as local details, global structures, or semantic relationships. They implement consistency constraints to minimize the differences in feature representations among different expert networks, encouraging similar outputs when processing the same input, thereby enhancing overall performance and robustness. The core idea of these methods is to design appropriate consistency constraints so that multi-branch or multi-expert networks can learn from each other during training, reducing their respective biases.
3.3.3 Multi-task Joint Loss
These works aim to enhance the effectiveness of multimodal compositional learning through carefully designed multi-task joint loss functions. They typically incorporate some form of knowledge distillation, contrastive learning, or adversarial learning, along with task-specific loss functions such as classification loss or triplet ranking loss.
The FAME-ViL and FashionViL models by Han et al. [60, 26] use distillation loss to transfer knowledge from single-task teacher models to multi-task student models, enhancing generalization and preventing negative transfer. FAME-ViL [60] combines distillation with mutual information maximization, guiding learning with teacher model outputs and improving parameter efficiency. FashionViL [26] uses contrastive learning, classification loss, and distillation loss, combining these to boost task-specific performance and foster synergy across tasks, improving retrieval accuracy and robustness.
Some works introduce adversarial learning mechanisms, leveraging generative adversarial networks to capture finer-grained semantic information. The multi-order adversarial network by Fu et al. [59] introduces an adversarial learning mechanism. In addition to adversarial losses, MAN combines triplet ranking loss to enhance the retrieval network’s performance, ensuring that the synthesized features better reflect the semantic information of text modifications. This comprehensive loss design effectively integrates adversarial learning with triplet loss, boosting the model’s performance in multimodal compositional retrieval tasks. Similarly, VISTA by Zhou et al. [10] combines cross-modal contrastive learning and bidirectional generative adversarial networks, aligning semantic information across modalities through adversarial training, improving both feature consistency and generation quality.
4 Zero-Shot Learning-based CMR
Without relying on triplet annotations, thereby avoiding limitations caused by data, existing methods focus on simulating compositional retrieval based solely on image-text information, establishing a zero-shot learning mechanism. Popular paradigms can be divided into three categories: (1) using images as input, designing a learning framework for visual-to-language transformation to combine modified text with query target images at test time; (2) using text as input, simulating the visual-to-language transformation process, and replacing text input with visual image input at test time, which offers higher training efficiency; (3) introducing additional knowledge for enhancement. The following sections will introduce the progress in each of these areas.
4.1 Image-side Transformation
In existing zero-shot compositional multi-modal retrieval, the process of image-to-visual transformation can be categorized into explicit and implicit approaches. The explicit approach directly converts input images into explicit textual descriptions, typically using pre-trained captioning models due to their training-free nature, but it is limited by the performance of these captioning models. The implicit approach, on the other hand, transforms input images into implicit feature vectors, usually encoded in a word vector space. This method is learnable and more flexible, allowing for a more sophisticated design tailored to the specific characteristics of the task.
4.1.1 Explicit visual Transformation (Training-free)
Existing explicit visual transformation methods [67, 68, 69] utilize pre-trained visual-language models (VLMs) and large language models (LLMs) to achieve zero-shot composed image retrieval (ZS-CIR) in a training-free manner.
Karthik et al. [67] propose a modular framework CIReVL, which generates captions for reference images and uses LLMs to recompose these captions based on the textual target modifications. This approach avoids the need for additional adaptation resources, supports flexible model component replacements, and allows user-level interventions by refining captions post-hoc, achieving competitive or superior results on four CIR benchmarks. Sun et al. [68] introduce a two-stage ranking method to address modality gaps and ambiguous requirements from the reference image. The first stage converts composed image-text queries into text-only queries for global retrieval, while the second stage re-ranks top-K results by extracting and evaluating local attributes from the modified instructions. This method combines global and local information to outperform other training-free approaches on open-domain datasets. Yang et al. [69] develop LDRE, which employs dense caption generation for reference images and uses LLMs for divergent reasoning to create a range of potential target image captions. An ensemble strategy integrates these candidates to better align with the query intent, effectively handling complex or ambiguous queries. Despite their differences in visual transformation strategies, all methods rely on pre-trained models and language-driven reasoning to bridge the gap between modalities and enable effective, training-free composed retrieval.
4.1.2 Implicit visual Transformation
Existing works on implicit visual transformation for zero-shot composed image retrieval (ZS-CIR) can be categorized into three main groups: single pseudo-word methods [70, 16, 71, 72], multiple pseudo-word methods [73, 74], and other methods [75, 76].
Single pseudo-word methods focus on mapping an image into a single pseudo-word token for textual representation. Saito et al. [70] propose Pic2Word, which maps CLIP visual features to a single pseudo-word token via a cycle contrastive loss, showing the powerful generalization across various ZS-CIR tasks using only weakly-supervised or unsupervised data. Similarly, Baldrati et al. [16] introduce SEARLE, an optimization-based textual inversion to generate pseudoword tokens and use knowledge distillation to train a Textual Inversion Network. SEARLE achieves efficient inference and improves performance. Moreover, they introduce an open-domain benchmarking dataset, CIRCO, which is the first CIR dataset to contain multiple ground-truth labels for each query. Expanding SEARLE, the iSEARLE proposed by Agnolucci et al. [71] further improves the retrieval accuracy by improving robustness with Gaussian noise, adding regularization to ensure dense token representations, and incorporating hard negative sampling to capture fine-grained details. Lastly, Context-I2W [72] employs a context-dependent mapping strategy, combining an Intent View Selector for dynamic view selection and a Visual Target Extractor for contextually relevant pseudo-word tokens, achieving significant performance gains.
Multiple pseudo-word methods extend representation by mapping images to multiple pseudo-word tokens to better capture fine-grained details. Lin et al. [73] introduces fine-grained textual inversion network, FTI4CIR, where an image is represented by a subject-oriented pseudo-word token and several attribute-oriented pseudo-word tokens, aligned with real-word tokens using semantic regularization. This approach enhances the expressiveness of image content. Similarly, Du et al. [74] propose an asymmetric ZS-CIR framework ISA, which employs an adaptive token learner to map images into sentence-like representations with discriminative visual information, and combines global contrastive distillation and local alignment regularization to improve retrieval accuracy and efficiency, particularly in resource-constrained environments.
Some approaches focus on improving the integration of image and text representations or adapting pre-trained models for ZS-CIR. The Slerp proposed by Jang et al. [76] employs Spherical Linear Interpolation for direct merging of image and text representations and Text-Anchored Tuning (TAT) to refine the image encoder, narrowing the modality gap and preserving image representation integrity. Chen et al. [75] propose Masked Tuning to bridge the gap between pre-trained models and ZS-CIR tasks by masking image patches to generate triplets (masked image, text, image), and learning fine-grained text-guided modifications that capture more detailed information for improved retrieval.
While single pseudo-word methods prioritize efficient and robust representations, multiple pseudo-word methods enhance the granularity of visual-text alignment. Other methods innovate in merging modalities or adapting pre-trained models for ZS-CIR, showcasing diverse yet complementary strategies to advance the field.
4.2 Text-side Transformation
The ZS-CIR methods for image-side transformation rely on pre-defined simple prompts (e.g., “a photo of [\$]”), which leads to a very homogeneous input to text encoders and cannot fully learn the subtleties of natural language variations. This limitation becomes more significant especially when trying to scale up the model, as larger models require more diverse data to capture more complex semantic information.
Gu et al. [77] and Byun et al. [78] focus on optimizing text-side transformations for CIR by addressing limitations in text encoding and modality alignment. The LinCIR proposed by Gu et al. leverages a self-supervised mechanism called Self-Masking Projection (SMP) to reduce dependency on triplet datasets and minimize modality gaps. This involves generating augmented texts through keyword masking and adding noise to textual embeddings, ensuring consistency between original and augmented representations. In contrast, Byun et al. introduce RTD, a plug-and-play training scheme that enhances text encoders through target-anchored contrastive learning, refined batch sampling with hard negatives, and a concatenation scheme, effectively addressing task discrepancies without requiring additional fine-tuning data.
Thawakar et al. [79] and Li et al. [80] emphasize integrating multimodal contexts to improve CIR and related tasks. Thawakar et al. propose CoVR, which uses rich language descriptions to encode query-specific contexts and learns discriminative visual-textual embeddings for accurate video retrieval, achieving sota results. Li et al. focus on multimodal compositional learning (MCL) for large language models, introducing tasks like Multimodal-Context Captioning and Retrieval to guide frozen language models in synthesizing multimodal information. Both methods enhance alignment and retrieval accuracy by leveraging complementary multimodal features, with MCL targeting more complex multimodal contexts, such as text generation and scene understanding.
4.3 External Knowledge Assistance
Suo et al. [81] introduce the Knowledge-Enhanced Dual-stream ZS-CIR (KEDs) framework to address the limitation of existing methods that overlook fine-grained attribute information. While traditional approaches mainly focus on global visual features, KEDs enhances pseudo-word tokens by integrating external knowledge from a database, which captures detailed attributes like color, object count, and layout. The framework includes two key components: a Bi-modal Knowledge-guided Projection (BKP) network that enriches pseudo-word tokens by retrieving relevant images and captions, and a Dual-stream Semantic Alignment module that explicitly aligns these enriched tokens with textual semantics. This approach leverages external knowledge to significantly improve the accuracy and domain-specific performance of ZS-CIR.
5 Semi-Supervised Learning-based CMR
In response to the problem that CIR datasets are small and difficult to scale, some works focus on semi-supervised learning-based Cross-Modal Retrieval (CMR), which can be broadly classified into four categories: (1) Automated data construction [1, 7]; (2) Synthetic data generation and dataset expansion [9, 5]; (3) Few-shot and semi-supervised learning [2, 6]; (4) Multimodal representation and retrieval enhancement [10, 8, 80]. Below, we will introduce the details of each category.
(1) Automated data construction: Liu et al. [1] introduce an automated data construction pipeline that leverages large-scale image-caption data to build training datasets for CIR task. The proposed TransAgg utilizes a transformer-based aggregation technique to dynamically fuse information from different modalities, demonstrating strong performance, particularly in zero-shot settings. Similarly, Levy et al. [7] focuse on overcoming the limitations of small datasets and modality redundancy in CIR systems. They introduce the LaSCo dataset, which is ten times larger than previous ones, and propose a CASE model, which integrates early fusion of modalities through a cross-attention mechanism. CASE improves retrieval performance on standard test sets. Both approaches highlight the critical role of large-scale, high-quality datasets and efficient fusion methods for enhancing CMR performance.
(2) Synthetic data generation and dataset expansion: Gu et al.[9] propose CompoDiff, a diffusion-based model designed to generate synthetic data for ZS-CIR. By creating a large-scale synthetic dataset, SynthTriplets18M, they extend the available training data, which significantly enhance the generalization ability of CIR models. CompoDiff also offers flexibility by handling conditions beyond traditional text, such as negative text and image mask conditions. Later, Jiang et al. [5] develope HyCIR, which uses synthetic labels generated from unlabeled images via a synthesis pipeline. This pipeline utilizes VLMs and LLMs to generate query text for each image pair, followed by semantic filtering. HyCIR employs a hybrid training strategy that combines zero-shot CIR supervision with synthetic triples to improve the generalization ability.
(3) Few-shot and semi-supervised learning: Hou et al. [2] address the challenge with limited labeled data by introducing a two-stage pseudo triplet guided few-shot CIR scheme (PTG-FSCIR). The first stage employs a masking training strategy and advanced image caption generator to construct pseudo triplets from pure image data; the second stage selects challenging unlabeled samples for fine-tuning based on active learning. This method demonstrates that carefully selected pseudo-data can significantly improve model performance, even with a small number of annotated samples. Jang et al. [6] propose Visual Delta Generator (VDG), a semi-supervised method that generates pseudo-triplets by describing visual differences between reference and target images using LLMs. The advantage of this method is that it can fully utilize existing large-scale image-text pair datasets without additional annotation costs.
(4) Multimodal representation and retrieval enhancement: Zhou et al. [10] propose VISTA, an embedding model for universal multimodal retrieval that overcomes the limitations of text-oriented retrievers by handling image-only and text-only data. VISTA generates high-quality composed image-text pairs using a template-based method and an LLM-based approach for diverse image descriptions. Then, VISTA employs a multi-stage training algorithm, aligning visual and textual encoders with weakly labeled data, and enhancing multimodal representation capability. Zhang et al. [8] introduce the CTI-IR model, which combines GANs and automatically manipulates the visual features of the reference image to match the textual description by jointly training the generative and retrieval models. Li et al. [80] introduce Multimodal Composition Learning (MCL), designed to improve LLMs’ understanding of complex visual and language contexts. MCL includes two tasks—Multimodal-Context Captioning and Multimodal-Context Retrieval—that enhance LLMs’ ability to process multimodal data. MCL shows strong performance on various benchmarks, especially in the zero-shot setting of the CIR task, revealing the potential of task-specific guidance for LLMs.
By leveraging techniques like automated data construction, synthetic data generation, few-shot learning, and advanced multimodal representations, these methods allow models to perform well and achieve high accuracy, even with limited data. These advancements improve model robustness and offer new solutions to the data bottleneck in CIR tasks.
6 Applications
At the application level, we categorize and summarize existing multimodal retrieval applications based on specific scenarios, including fashion images, natural images, videos, remote sensing images, pedestrian re-identification, and skeletal diagrams. The specific contexts and challenges faced by each task in these different applications are thoroughly summarized, with the hope of inspiring exploration in other yet-to-be-explored scenarios.
6.1 Composed Fashion Image Retrieval
Traditional fashion image retrieval primarily relies on simple image search or keyword-based search. However, these methods often fail to meet user needs when searching for specific fashion items with complex attributes such as color, style, and material. The main goal of Composed Fashion Image Retrieval (CFIR) is to achieve more accurate and personalized fashion searches by combining both images and texts. Users can input a reference image (e.g., a photo of a clothing item) along with a textual modification (e.g., ”long-sleeve version”), allowing the system to understand both the visual and semantic information, leading to more refined retrieval.
CFIR can be widely applied on e-commerce platforms, helping users quickly find products that meet specific needs, thus improving the convenience and accuracy of the purchase. Users can not only search by image but also refine their queries with textual descriptions, improving the efficiency and quality of search results. This increases the precision of online shopping, making it easier for users to find products that match their requirements, boosting customer satisfaction, and increasing overall user engagement. In addition, for retailers and brands, CFIR enables more accurate product recommendations based on user preferences while also supporting fashion trend analysis and inventory management.
6.1.1 Benchmark Datasets
The commonly used datasets related to the retrieval of composed fashion images can be summarized as follows:
FashionIQ [82] consists of 77,684 images of fashion products across three categories: dresses, shirts, and tops&tees. A subset of 49,464 images is annotated with side information derived from product descriptions, including various attributes. Additionally, 60,272 image pairs are annotated with relative captions, which provide natural language descriptions of the differences between reference and target images. These captions serve as modification text for retrieving target images.
Fashion200k [83] is collected from online shopping websites. The authors crawl over 300,000 product images along with their descriptions, removing those with descriptions containing fewer than four words, resulting in more than 200,000 images. They divide the dataset into 172,049 images for training, 12,164 for validation, and 25,331 for testing. For text cleaning, they eliminate stop words, symbols, and words that occur fewer than five times. The remaining words are treated as attributes, giving a total of 4,404 attributes to train the joint embedding.
Shoes [84] is collected from the Internet, specifically from like.com, a shopping website that aggregates product data from a wide range of e-Commerce sources. They crawl 10,751 pairs of shoe images with relative expressions that describe fine-grained visual differences. Typically, 10,000 samples are used for training and 4,658 samples are used for evaluation.
CSS [18] consists of 32K queries (16K for training and 16K for test), such as ’make yellow sphere small’ that serve as modification text for the images synthesized in a 3-by-3 grid scene. Although it is a relatively simple dataset, CSS has the benefit of facilitating carefully controlled experiments.
CFQ [17] is collected from a set of images from online advertisements from various fashion brands and retailers that feature one dress, usually worn by a model. It has explicit positive and negative labels for all images in the same category as responses to each query. The data set contains 334 images with 3 categories of dress, dress, and sundress and 40 queries.
6.1.2 Results and Analysis
TABLE I: Comparison of Different Methods on FashionIQ Dataset(Original Split)
Method Backbone Dresses Shirts Tops&Tees Avg
Visual Textual R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
Supervised Learning-based CMR (SL-CMR)
JSVM20’ECCV MobileNet LSTM $10.70$ $25.90$ $12.00$ $27.10$ $13.00$ $26.90$ $11.90$ $26.63$
TRACE21’AAAI ResNet50 GRU $22.70$ $44.91$ $20.80$ $40.80$ $24.22$ $49.80$ $22.57$ $46.19$
ComposeAE21’WACV ResNet18 BERT $10.77$ $28.29$ $9.96$ $25.14$ $12.74$ $30.79$ $11.80$ $29.40$
CoSMo21’CVPR ResNet50 LSTM $21.39$ $44.45$ $16.90$ $37.49$ $21.32$ $46.02$ $19.87$ $42.62$
FashionVLP22’CVPR ResNet50 BERT $26.77$ $53.2$ $22.67$ $46.22$ $28.51$ $57.47$ $25.98$ $52.30$
Combiner22’CVPR-D CLIP(RN50) CLIP $31.63$ $56.67$ $36.36$ $51.86$ $38.19$ $62.42$ $35.39$ $59.03$
MultiColSAP22’TMM ResNet50 LSTM $30.74$ $55.92$ $26.05$ $50.64$ $34.42$ $61.14$ $30.4$ $55.9$
CLIP4Cir122’CVPR-W CLIP(RN50) CLIP $33.81$ $59.40$ $39.99$ $60.45$ $41.41$ $65.37$ $38.32$ $61.74$
Progressive22’SIGIR ResNet50 BERT $29.00$ $53.94$ $35.43$ $58.88$ $39.16$ $64.56$ $34.53$ $59.13$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $33.60$ $58.90$ $39.45$ $61.78$ $43.96$ $68.33$ $39.02$ $63.00$
ARTEMIS22’ICLR ResNet50 BiGRU $25.68$ $51.05$ $21.57$ $44.13$ $28.59$ $55.06$ $25.28$ $50.08$
ComqueryFormer23’TMM Swin-T BERT $28.85$ $55.38$ $25.64$ $50.22$ $33.61$ $60.48$ $29.37$ $55.36$
TransAgg(LaionCIR)23’BMVC BLIP BLIP $27.67$ $49.38$ $32.83$ $52.31$ $35.70$ $58.08$ $32.07$ $53.26$
CLIP-CD23’AI CLIP(RN50) CLIP $37.68$ $62.62$ $42.44$ $63.74$ $45.33$ $67.72$ $41.82$ $64.79$
CLIP4Cir223’TOMM CLIP(RN50) CLIP $37.67$ $63.16$ $39.87$ $60.84$ $44.88$ $68.59$ $40.80$ $64.20$
BLIP4CIR24’WACV BLIP BLIP $40.65$ $66.34$ $37.49$ $60.06$ $43.60$ $67.77$ $39.34$ $63.85$
BLIP4CIR+Bi24’WACV BLIP BLIP $42.09$ $67.33$ $41.76$ $64.28$ $46.61$ $70.32$ $43.49$ $67.31$
SPIRIT24’TOMM CLIP(RN50) CLIP $39.86$ $64.30$ $47.67$ $71.10$ $44.11$ $65.60$ $43.88$ $67.20$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $51.90$ $74.37$ $53.57$ $73.21$ $58.48$ $79.23$ $54.65$ $75.60$
FashionViL24’ECCV ResNet50 BERT $33.47$ $59.94$ $25.17$ $50.39$ $34.98$ $60.79$ $31.21$ $57.04$
BLIP4CIR224’TMLR BLIP(ViT-B) BLIP $43.78$ $71.34$ $50.15$ $71.25$ $55.23$ $76.80$ $51.17$ $73.13$
CompoDiff23’TMLR CLIP(ViT-L/14) $35.53$ $49.56$ $40.88$ $53.06$ $41.15$ $54.12$ $39.05$ $52.34$
CompoDiff23’TMLR CLIP(ViT-G/14) $38.39$ $51.03$ $41.68$ $56.02$ $45.70$ $57.32$ $39.81$ $51.90$
SDQUR24’TCSVT BLIP2 $49.93$ $73.33$ $56.87$ $76.50$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM CLIP LSTM $35.75$ $60.56$ $37.02$ $60.05$ $42.25$ $67.52$ $38.20$ $62.82$
NSFSE24’TMM ResNet50 BiGRU $29.62$ $54.41$ $22.96$ $45.93$ $31.08$ $57.01$ $27.84$ $52.39$
NSFSE24’TMM ResNet152 BiGRU $31.12$ $55.73$ $24.58$ $45.85$ $31.93$ $58.37$ $29.17$ $53.24$
MANME24’TCSVT ResNet50 BiGRU $31.26$ $57.66$ $26.37$ $47.94$ $36.33$ $59.31$ $29.95$ $54.90$
Zero-Shot Learning-based CMR (ZSL-CMR)
PALAVAR22’ECCV CLIP(ViT-B/32) $17.25$ $35.94$ $21.49$ $37.05$ $20.55$ $38.76$ $19.76$ $37.25$
Pic2Word23’CVPR CLIP(ViT-L/14) $20.00$ $40.20$ $26.20$ $43.60$ $27.90$ $47.40$ $24.70$ $43.70$
SEARLE23’ICCV CLIP(ViT-B/16) $18.54$ $38.51$ $24.44$ $41.61$ $25.70$ $46.46$ $22.89$ $42.53$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $21.57$ $44.47$ $30.37$ $47.49$ $30.9$ $51.76$ $27.61$ $47.9$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $20.48$ $43.13$ $26.86$ $45.58$ $29.32$ $49.97$ $25.56$ $46.23$
MTCIR23’Arxiv CLIP(ViT-L/14) $28.11$ $51.12$ $38.63$ $58.51$ $39.42$ $62.68$ $35.39$ $57.44$
VQA4CIR(+CLIP4CIR)23’Arxiv LLaMA + LLaVA $40.91$ $65.13$ $45.62$ $65.68$ $49.21$ $71.22$ $45.24$ $67.34$
VQA4CIR(+SPRC)23’Arxiv LLaMA + LLaVA $49.18$ $73.06$ $56.79$ $74.52$ $59.67$ $79.30$ $55.21$ $75.62$
KEDs24’CVPR CLIP(ViT-L/14) $21.7$ $43.8$ $28.9$ $48.0$ $29.9$ $51.9$ $26.8$ $47.9$
PM24’Arxiv CLIP(ViT-L/14) $27.1$ $43.8$ $21.4$ $41.7$ $28.9$ $47.3$ $25.8$ $44.2$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $24.39$ $47.84$ $31.35$ $50.59$ $32.43$ $54.21$ $29.39$ $50.88$
CIReVL24’ICLR CLIP(ViT-B/32) $25.29$ $46.36$ $28.36$ $47.84$ $31.21$ $53.85$ $28.29$ $49.35$
CIReVL24’ICLR CLIP(ViT-L/14) $24.79$ $44.76$ $29.49$ $47.40$ $31.36$ $53.65$ $28.55$ $48.57$
CIReVL24’ICLR CLIP(ViT-G/14) $27.07$ $49.53$ $33.71$ $51.42$ $35.80$ $56.14$ $32.19$ $52.36$
LinCIR24’CVPR CLIP(ViT-L/14) $20.92$ $42.44$ $29.10$ $46.81$ $28.81$ $50.18$ $26.28$ $46.49$
LinCIR24’CVPR CLIP(ViT-H/14) $29.80$ $52.11$ $36.90$ $57.75$ $42.07$ $62.52$ $36.26$ $57.46$
LinCIR24’CVPR CLIP(ViT-G/14) $38.08$ $60.88$ $46.76$ $65.11$ $50.48$ $71.09$ $45.11$ $65.69$
iSEARLE-XL-OTI24’Arxiv CLIP(ViT-L/14) $24.19$ $45.12$ $31.8$ $50.2$ $31.72$ $53.29$ $29.24$ $49.54$
iSEARLE-XL24’Arxiv CLIP(ViT-L/14) $22.51$ $46.36$ $28.75$ $47.84$ $31.31$ $52.68$ $27.52$ $48.96$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $20.72$ $43.13$ $26.69$ $44.31$ $26.67$ $48.75$ $24.7$ $45.4$
RTD((LinCIR)24’Arxiv CLIP(ViT-B/32) $19.98$ $41.75$ $23.65$ $42.74$ $24.73$ $46.56$ $22.79$ $43.68$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $23.2$ $47.25$ $32.63$ $50.39$ $32.18$ $54.56$ $29.34$ $50.73$
RTD((LinCIR)24’Arxiv CLIP(ViT-L/14) $24.49$ $48.24$ $32.83$ $50.44$ $33.4$ $54.56$ $30.24$ $51.08$
ISA(Sym)24’ICLR BLIP $24.69$ $43.88$ $30.79$ $50.05$ $33.91$ $53.65$ $29.79$ $49.19$
ISA(Asy)24’ICLR EfficientNet+BLIP $25.33$ $46.26$ $30.03$ $48.58$ $33.45$ $53.8$ $29.6$ $49.54$
LDRE24’SIGIR CLIP(ViT-B/32) $19.97$ $41.84$ $27.38$ $46.27$ $27.07$ $48.78$ $24.81$ $45.63$
LDRE24’SIGIR CLIP(ViT-L/14) $22.93$ $46.76$ $31.04$ $51.22$ $31.57$ $53.64$ $28.51$ $50.54$
LDRE24’SIGIR CLIP(ViT-G/14) $26.11$ $51.12$ $35.94$ $58.58$ $35.42$ $56.67$ $32.49$ $55.46$
Context-I2W24’AAAI CLIP(ViT-L/14) $23.1$ $45.3$ $29.7$ $48.6$ $30.6$ $52.9$ $27.8$ $48.9$
Slerp24’ECCV CLIP(ViT-B/32) $20.53$ $41.00$ $23.75$ $40.92$ $26.98$ $46.77$ $23.75$ $42.90$
Slerp+TAT24’ECCV CLIP(ViT-B/32) $19.24$ $42.14$ $23.06$ $41.95$ $26.57$ $47.78$ $22.96$ $43.96$
Slerp+TAT24’ECCV CLIP(ViT-L/14) $23.35$ $45.12$ $29.94$ $46.47$ $31.97$ $51.20$ $28.32$ $47.00$
Slerp+TAT24’ECCV BLIP(ViT-L/16) $29.15$ $50.62$ $32.14$ $51.62$ $37.02$ $57.73$ $32.77$ $53.32$
HyCIR(CC3M+syn)24’Arxiv CLIP $19.98$ $40.8$ $27.62$ $44.94$ $28.14$ $47.67$ $25.25$ $44.47$
HyCIR(CC3M+syn)24’Arxiv BLIP $18.88$ $34.5$ $22.52$ $37.58$ $22.13$ $40.33$ $21.18$ $37.47$
Semi-Supervised Learning-based CMR (SSL-CMR)
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $32.24$ $46.27$ $37.69$ $49.08$ $38.12$ $50.57$ $36.02$ $48.64$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $37.78$ $49.10$ $41.31$ $55.17$ $44.26$ $56.41$ $39.02$ $51.71$
PTG(+SPRC)24’Arxiv $31.10$ $51.90$
CASE24’AAAI BLIP $47.44$ $69.36$ $48.48$ $70.23$ $50.18$ $72.24$ $48.79$ $70.68$
VDG(+DF’se)24’CVPR BLIP $47.10$ $69.10$ $49.95$ $69.96$ $53.90$ $74.35$ $50.32$ $71.14$
VDG(+FIQ’se)24’CVPR BLIP $47.89$ $69.81$ $51.36$ $71.08$ $53.29$ $74.65$ $50.85$ $71.85$
TABLE II: Comparison of Different Methods on FashionIQ Dataset(Val Split)
Method Backbone Dresses Shirts Tops&Tees Avg
Visual Textual R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
Supervised Learning-based CMR (SL-CMR)
TIRG18’CVPR ResNet18 LSTM $14.87$ $34.66$ $18.26$ $37.89$ $19.08$ $39.62$ $17.40$ $37.39$
VAL20’CVPR ResNet50 LSTM $22.53$ $44.00$ $22.38$ $44.15$ $27.53$ $51.68$ $24.15$ $46.61$
MAAF20’Arxiv ResNet50 LSTM $23.8$ $48.6$ $21.3$ $44.2$ $27.9$ $53.6$ $24.3$ $48.8$
TRACE21’AAAI ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.7$ $61.23$ $29.08$ $54.7$
RTIC21’Arxiv ResNet50 LSTM $19.4$ $43.51$ $16.93$ $38.36$ $21.58$ $47.88$ $19.3$ $43.25$
DCNet21’AAAI ResNet50 Glove $28.95$ $56.07$ $23.95$ $47.30$ $30.44$ $58.29$ $27.78$ $53.89$
CoSMo21’CVPR ResNet50 LSTM $25.64$ $50.3$ $24.90$ $49.18$ $29.21$ $57.46$ $26.58$ $52.31$
HFCA21’ACMMM ResNet50 LSTM $26.2$ $51.2$ $22.4$ $46.0$ $29.7$ $56.4$ $26.1$ $51.2$
CLVC-Net21’SIGIR ResNet50 LSTM $29.85$ $56.47$ $28.75$ $54.76$ $33.50$ $64.00$ $30.70$ $58.41$
CIRPLANT21’ICCV ResNet152 LSTM $14.38$ $34.66$ $13.64$ $33.56$ $16.44$ $38.34$ $14.82$ $35.52$
CIRPLANT W/ OSCAR21’ICCV ResNet152 LSTM $17.45$ $40.41$ $17.53$ $38.81$ $21.64$ $45.38$ $18.87$ $41.53$
JPM(VAL+MSE)21’ACMMM ResNet18 LSTM $21.27$ $43.12$ $21.88$ $43.3$ $25.81$ $50.27$ $22.98$ $45.59$
JPM(VAL+Tri)21’ACMMM ResNet18 LSTM $21.38$ $45.15$ $22.81$ $45.18$ $27.78$ $51.70$ $23.99$ $47.34$
FashionVLP22’CVPR ResNet50 BERT $32.42$ $60.29$ $31.89$ $58.44$ $38.51$ $68.79$ $34.27$ $62.51$
SAC22’WACV ResNet50 GRU $26.52$ $51.01$ $28.02$ $51.86$ $32.70$ $61.23$ $29.08$ $54.70$
ERR22’TIP ResNet50 LSTM $30.02$ $55.44$ $25.32$ $49.87$ $33.20$ $60.34$ $29.51$ $55.22$
CRR22’ACMMM ResNet101 BiGRU $30.41$ $57.11$ $30.73$ $58.02$ $33.67$ $64.48$ $31.60$ $59.87$
Progressive22’SIGIR ResNet50 BERT $33.22$ $59.99$ $46.17$ $68.79$ $46.46$ $73.84$ $41.98$ $67.54$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $38.18$ $64.50$ $48.63$ $71.54$ $52.32$ $76.90$ $46.37$ $70.98$
MCR22’TMM ResNet50 LSTM $26.20$ $51.20$ $22.40$ $46.00$ $29.70$ $56.40$ $26.10$ $51.20$
NEUCORE23’NIPS-W ResNet BiGRU $27.00$ $53.79$ $22.84$ $45.00$ $29.63$ $56.65$ $26.45$ $51.75$
ARTEMIS22’ICLR ResNet50 LSTM $27.34$ $51.71$ $21.05$ $49.87$ $27.34$ $44.18$ $24.43$ $48.59$
ARTEMIS22’ICLR ResNet50 BiGRU $27.16$ $52.40$ $21.78$ $43.64$ $29.20$ $54.83$ $26.05$ $50.29$
AMC23’TOMM ResNet50 LSTM $31.73$ $59.25$ $36.21$ $66.60$ $30.67$ $59.08$ $32.87$ $61.64$
ComqueryFormer23’TMM Swin-T BERT $33.86$ $61.08$ $35.57$ $62.19$ $42.07$ $69.30$ $37.17$ $64.19$
RankUn23’Arxiv CLIP(RN50) CLIP $34.8$ $60.22$ $45.21$ $69.06$ $47.68$ $74.85$ $42.5$ $68.04$
CRN23’TIP Swin-T(base) LSTM $30.24$ $57.61$ $29.83$ $55.54$ $33.91$ $64.04$ $31.36$ $59.06$
CRN23’TIP Swin-T(Large) LSTM $32.67$ $59.30$ $30.27$ $56.97$ $37.74$ $65.94$ $33.56$ $60.74$
SSN23’Arxiv CLIP(ViT-B/32) CLIP $34.36$ $60.78$ $38.13$ $61.83$ $44.26$ $69.025$ $38.92$ $63.89$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $45.22$ $69.66$ $52.60$ $72.52$ $56.14$ $77.10$ $51.32$ $73.09$
MCEMiner24’TIP ResNet50 LSTM $33.23$ $59.16$ $26.15$ $50.87$ $33.83$ $61.40$ $31.07$ $57.14$
CaLa-CLIP4Cir24’SIGIR $32.93$ $56.82$ $39.20$ $60.13$ $39.16$ $63.83$ $37.10$ $60.26$
CaLa-ARTEMIS24’SIGIR $40.13$ $66.88$ $46.86$ $67.28$ $49.87$ $74.11$ $45.62$ $69.42$
CaLa-BLIP2cir24’SIGIR $42.38$ $66.08$ $46.76$ $68.16$ $50.93$ $73.42$ $46.69$ $69.22$
CSS-Net24’KBS ResNet50 RoBERTa $33.65$ $63.16$ $35.96$ $61.96$ $42.65$ $70.70$ $37.42$ $65.27$
SPIRIT24’TOMM CLIP(RN50) CLIP $43.83$ $68.86$ $52.50$ $74.19$ $56.60$ $79.25$ $50.98$ $74.10$
LIMN24’TPAMI ResNet50 LSTM $35.6$ $62.37$ $34.69$ $59.81$ $40.64$ $68.33$ $36.98$ $63.50$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $57.63$ $78.56$ $62.14$ $80.38$ $66.15$ $85.73$ $61.97$ $81.56$
Uncertainty24’ICLR ResNet50 RoBERTa $30.60$ $57.46$ $31.54$ $58.29$ $37.37$ $68.41$ $33.17$ $61.39$
Uncertainty(CLVC-Net)24’ICLR ResNet50 RoBERTa $31.25$ $58.35$ $31.69$ $60.65$ $39.82$ $71.07$ $34.25$ $63.36$
Uncertainty(CLIP4Cir)24’ICLR ResNet50 RoBERTa $32.61$ $61.34$ $33.23$ $62.55$ $41.40$ $72.51$ $35.75$ $65.47$
SDQUR24’TCSVT BLIP2 $56.87$ $76.50$ $49.93$ $73.33$ $59.66$ $79.25$ $55.49$ $76.36$
AlRet24’TMM ResNet50 LSTM $30.19$ $58.80$ $29.39$ $55.69$ $37.66$ $64.97$ $32.26$ $59.76$
SPRC24’ICLR BLIP2(ViT-L) $49.18$ $72.43$ $55.64$ $73.89$ $59.35$ $78.58$ $54.92$ $74.97$
Zero-Shot Learning-based CMR (ZSL-CMR)
Progressive(stage1)22’SIGIR ResNet50 BERT $5.75$ $13.04$ $11.73$ $21.98$ $11.22$ $21.93$ $9.57$ $18.98$
TABLE III: Comparison of Different Methods on Fashion200K & Shoes Datasets
Method Backbone Fashion200K Shoes
Visual Textual R@1 R@10 R@50 Avg R@1 R@10 R@50 Avg
Supervised Learning-based CMR (SL-CMR)
TIRG18’CVPR ResNet18 LSTM $14.1$ $42.5$ $63.8$ $40.1$
VAL20’CVPR MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
LBF(small)20’CVPR Faster-RCNN $16.26$ $46.9$ $71.73$ $44.96$
LBF(big)20’CVPR Faster-RCNN $17.8$ $48.35$ $68.5$ $44.83$
VAL20’CVPR MobileNet&ResNet50 LSTM $22.9$ $50.8$ $72.7$ $48.8$ $17.18$ $51.52$ $75.83$ $48.18$
JSVM20’ECCV MobileNet LSTM $19.0$ $52.1$ $70.0$ $47.0$
TRACE21’AAAI ResNet50 GRU $18.5$ $51.73$ $77.28$
MAAF20’Arxiv ResNet50 LSTM $18.94$
GSCMR21’TIP Faster-RCNN BiGRU $21.57$ $52.84$ $70.12$ $48.18$
MAN21’ICASSP ResNet18 BERT $17.1$ $47.9$ $68.1$ $44.37$
MAN21’ICASSP MobileNet BERT $22.3$ $54.5$ $74.1$ $50.3$
RTIC21’Arxiv ResNet50 LSTM $43.66$ $72.11$
ComposeAE21’WACV ResNet18 BERT $22.8$ $55.3$ $73.4$ $50.5$
DCNet21’AAAI ResNet50 Glove $46.89$ $67.56$ $53.82$ $79.33$
CoSMo21’CVPR ResNet18&ResNet50 LSTM $23.3$ $50.4$ $69.3$ $47.7$ $16.72$ $48.36$ $75.64$ $46.91$
HFCA21’ACMMM ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$ $17.85$ $50.95$ $77.24$ $48.68$
CLVC-Net21’SIGIR ResNet50 LSTM $22.6$ $53.0$ $72.2$ $49.3$ $17.64$ $54.39$ $79.47$ $50.50$
HCL21’MMAsia ResNet18 LSTM $23.48$ $54.03$ $73.71$ $50.41$
JPM(TIRG+MSE)21’ACMMM ResNet18 LSTM $19.8$ $46.5$ $66.6$ $44.3$
JPM(TIRG+Tri)21’ACMMM ResNet18 LSTM $17.7$ $44.7$ $64.5$ $42.3$
FashionVLP22’CVPR ResNet18&ResNet50 BERT $49.9$ $70.5$ $60.2$ $49.08$ $77.32$ $63.20$ $63.20$
SAC22’WACV ResNet50 GRU $18.5$ $51.73$ $77.28$ $49.17$
ERR22’TIP ResNet50 LSTM $50.88$ $70.60$ $19.87$ $55.96$ $79.58$ $51.80$
TriArea22’Sci.Rep. ResNet18+TF LSTM $17.7$ $46.8$ $66.2$ $43.6$
CRR22’ACMMM ResNet101 GRU&BiGRU $24.85$ $56.41$ $73.56$ $51.61$ $19.41$ $56.38$ $79.92$ $51.90$
MultiColSAP22’TMM ResNet50 LSTM $51.06$ $70.13$ $60.60$ $19.54$ $65.39$ $79.47$ $54.80$
MCR22’TMM ResNet50 LSTM $18.24$ $49.41$ $69.37$ $45.67$
Progressive22’SIGIR ResNet50 BERT $19.53$ $55.65$ $80.58$ $51.92$
Progressive22’SIGIR CLIP(ViT-B/32) BERT $22.88$ $58.53$ $84.16$ $55.29$
ComqueryFormer23’TMM Swin-T BERT $52.2$ $72.2$ $62.2$
NEUCORE23’NIPS-W ResNet BiGRU $19.76$ $55.48$ $80.75$
ARTEMIS22’ICLR ResNet50 LSTM $17.60$ $51.05$ $76.85$ $48.50$
ARTEMIS22’ICLR ResNet50 BiGRU $18.72$ $53.11$ $79.31$ $50.38$
AMC23’TOMM ResNet50 LSTM $19.99$ $56.89$ $79.27$
CRN23’TIP Swin-T(base) LSTM $17.32$ $54.15$ $79.34$ $50.27$ $53.3$ $73.3$ $63.3$ $63.3$
CRN23’TIP Swin-T(Large) LSTM $18.92$ $54.55$ $80.04$ $51.17$ $53.5$ $74.5$ $64.0$ $64.0$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $58.05$
MCEMiner24’TIP ResNet18&ResNet50 LSTM $26.82$ $56.76$ $76.91$ $53.50$ $19.10$ $55.37$ $79.57$ $51.35$
CSS-Net24’KBS RoBERTa RoBERTa $22.2$ $50.5$ $69.7$ $47.5$ $20.13$ $56.81$ $81.32$ $52.75$
SPIRIT24’TOMM CLIP(RN50) CLIP $55.2$ $73.6$ $56.90$ $81.49$ $69.19$ $69.19$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $36.8$ $67.9$ $87.8$ $64.1$ $31.47$ $69.19$ $88.52$ $63.06$
LIMN24’TPAMI ResNet50 LSTM $57.30$ $82.7$
Uncertainty24’ICLR ResNet50 RoBERTa $21.8$ $52.1$ $70.2$ $48.0$ $18.41$ $53.63$ $79.84$ $50.63$
SDQUR24’TCSVT BLIP2 $30.14$ $68.30$ $88.30$ $62.25$
AlRet24’TMM ResNet18&ResNet50 LSTM $24.42$ $53.93$ $73.25$ $53.53$ $18.13$ $53.98$ $78.81$ $50.31$
AlRet24’TMM CLIP LSTM $21.02$ $55.72$ $80.77$ $52.50$
NSFSE24’TMM ResNet50 BiGRU $25.3$ $53.8$ $73.5$ $50.87$
NSFSE24’TMM ResNet152 BiGRU $24.9$ $54.3$ $73.4$ $50.87$
MANME24’TCSVT ResNet50 BiGRU $23.0$ $57.9$ $75.3$ $52.0$ $20.73$ $55.96$ $80.98$ $52.56$
TABLE IV: Comparison of Different Methods on CSS Dataset
Methods Backbone Recall@1
Visual Textual 3D-to-3D 2D-to-3D
Supervised Learning-based CMR (SL-CMR)
TIRG18’CVPR ResNet18 LSTM $73.70$ $46.60$
LBF(small)20’CVPR Faster-RCNN $67.26$ $50.31$
LBF(big)20’CVPR Faster-RCNN $79.20$ $55.69$
MAAF20’Arxiv ResNet50 LSTM $87.80$
GSCMR21’TIP Faster-RCNN BiGRU $81.81$ $58.74$
MAN21’ICASSP ResNet18 BERT $79.60$
MAN21’ICASSP MobileNet BERT $80.40$
HCL21’MMAsia ResNet18 LSTM $81.59$ $58.65$
JPM(TIRG+MSE)21’ACMMM ResNet18 LSTM $83.80$
JPM(TIRG+Tri)21’ACMMM ResNet18 LSTM $83.20$
CRR22’ACMMM ResNet101 BiGRU $85.84$
6.2 Composed Natural Image Retrieval
Traditional image retrieval systems often rely on single visual or textual information to retrieve the target image. However, this way is limited by the representation ability of a single modality, which makes it difficult to deal with complex image retrieval problems that involve multiple attributes or concepts. The main goal of Composed Natural Image Retrieval (CNIR) is to enhance the retrieval capability by composing the textual description with the image content, enabling the system to understand and respond to more abstract and complex user requirements. For instance, users can upload a photo of their favorite landscape with a description (e.g., “same location but in the Autumn”), and the CNIR system will parse the content of the image with the textual description, and understand the semantic associations to return a matched image. This technological advancement allows image retrieval to go beyond content similarity and capture deeper conceptual compositions.
CNIR is widely applied in various real-world scenarios, which has significant application value and business potential. In the field of travel, users can quickly find other images of the same scenic spot in a specific environment by uploading photos of the desired location and adding conditions such as season, time, and weather, providing inspiration and practical information for travel planning. Scientists can use the technology to conduct interdisciplinary research, such as ecologists combining geographic location and species image features to search for images in relevant literature. Furthermore, in the field of cultural heritage preservation, artists can use CNIR to compare images of artifacts from different periods of history and assist in restoration works. This precise and personalized search not only improves search efficiency and accuracy but also increases user participation and satisfaction, bringing new service models and development opportunities for the wide application of retrieval tasks.
6.2.1 Benchmark Datasets
CIRR [42] is the first released dataset for the natural domain. It consists of a total of 36554 triples derived from 21552 real-life images from the popular natural language reasoning dataset NLVR2 [85]. Each triplet consists of real-life images and human-generated modified sentences, arranged in an 80%-10%-10% split between the train/validate/test. This dataset encompasses rich object interactions, which addresses the issues of overly narrow domains and the high number of false negatives in the FashionIQ dataset [82]. Each query image in the validation and test sets has only one target image, where the test is evaluated on an online platform.
CIRCO [16] is the first dataset for CIR with multiple ground truths collected from the COCO 2017 unlabeled set [27], which is specifically designed for ZS-CIR evaluation. It consists of a total of 1020 queries, 220 for the validation set, and 800 for the test set, with an average of 4.53 ground truths per query. In addition, it uses all the 120K images of COCO as the index set, thus providing significantly more distractors than the 2K images of the CIRR test set. Compared to CIRR [42], CIRCO employs a two-phase annotation strategy to ensure higher quality, reduced false negatives, and the availability of multiple ground truths. Thus, it is also extremely challenging. The test is similarly evaluated on an online platform.
LaSCo [7] has 10 times more queries, 2 times more unique tokens, and 17 times more corpus images than the CIRR dataset [42]. It contains over 389K triplets of query image, modification text, and target image. The training image corpus has 81,653 images and the validation corpus has 39,826 images. The dataset has 13,488 unique language tokens and an average text length of 30.70 tokens per query. Analysis of the dataset shows it has significantly less bias towards a single modality for retrieval compared to previous datasets.
Laion-CIR-Combined [1] consists of two sub-datasets, Laion-CIR-Template and Laion-CIR-LLM, which are constructed by two methods of modifying captions (i.e., using templates or LLMs). Specifically, for one image-caption sample, they revise its caption and use the resulting edited caption as a query to retrieve an image with a similar caption as the target image. Both datasets contain around 16K triplets. In addition, by combining the two approaches, a 32K dataset.
SynthTriplets18M [9] is a vast set of high-quality 18M triplet datasets synthesized using large-scale generative models such as OPT [86] and Stable Diffusion [87, 88]. Specifically, they use two strategies: the keyword-based generation (11.4M), and the LLM-based generation (7.4M). SynthTriplets18M is over 500 times larger than existing datasets and covers a diverse and extensive range of conditioning cases.
Birds-to-Words [89] consists of images of birds from iNaturalist combined with human-annotated paragraphs to describe the difference between these pairs of images. This dataset is characterized by “long” natural language queries with every one of 3,347 image pairs having on average 4.8 paragraphs, each describing the differences between the pair of birds in an average of 31.38 words. Birds-to-Words provides richer text descriptions in each example than any of the other datasets current study, although the number of examples is small.
MIT-States [90] has 60k images, and each comes with an object/noun label and a state/adjective label (such as “red tomato” or “new camera”). For nouns, it selected words that refer to physical objects, materials, and scenes. For adjectives, it selected words that refer to specific physical transformations. Then, for each adjective, if a clear antonym exists, it is paired with another antonym adjective in the list. Finally, there are 249 nouns and 115 adjectives, on average each noun is only modified by 9 adjectives it affords.
6.2.2 Results and Analysis
TABLE V: Comparison of Different Methods on CIRR Dataset
Method Backbone Recall@K Recallsubset@K
Visual Textual K=1 K=5 K=10 K=1 K=2 K=3
Supervised Learning-based CMR (SL-CMR)
CIRPLANT21’ICCV ResNet152 LSTM $15.18$ $43.36$ $60.48$ $33.81$ $56.99$ $75.40$
CIRPLANT W/ OSCAR21’ICCV ResNet152 LSTM $19.55$ $52.55$ $68.39$ $39.20$ $63.03$ $79.49$
Combiner22’CVPR-D CLIP(RN50) CLIP $33.59$ $65.35$ $77.35$ $62.39$ $81.81$ $92.02$
CLIP4Cir122’CVPR-W CLIP(RN50) CLIP $38.53$ $69.98$ $81.86$ $68.19$ $85.64$ $94.17$
NEUCORE23’NIPS-W ResNet BiGRU $18.46$ $49.40$ $63.57$ $44.27$ $67.06$ $78.92$
ComqueryFormer23’TMM Swin-T BERT $25.76$ $61.76$ $75.90$ $51.86$ $76.26$ $89.25$
ARTEMIS22’ICLR ResNet50 BiGRU $16.96$ $46.10$ $61.31$ $39.99$ $62.20$ $75.67$
RankUn23’Arxiv CLIP(RN50) CLIP $32.24$ $66.63$ $79.23$ $61.25$ $81.33$ $92.02$
CLIP4Cir223’TOMM CLIP(RN50) CLIP $42.05$ $76.13$ $86.51$ $70.15$ $87.18$ $94.40$
SSN23’Arxiv CLIP(ViT-B/32) CLIP $43.91$ $77.25$ $86.48$ $71.76$ $88.63$ $95.54$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $45.25$ $78.29$ $87.16$ $72.84$ $89.25$ $94.13$
MCEMiner24’TIP ResNet152 LSTM $17.48$ $46.13$ $62.17$
CaLa-CLIP4Cir24’SIGIR $35.37$ $68.89$ $80.04$ $66.68$ $84.65$ $93.42$
CaLa-ARTEMIS24’SIGIR $47.37$ $79.33$ $88.17$ $76.02$ $90.29$ $96.19$
CaLa-BLIP2cir24’SIGIR $49.11$ $81.21$ $89.59$ $76.27$ $91.04$ $96.46$
BLIP4CIR24’WACV BLIP BLIP $40.17$ $71.81$ $83.18$ $72.34$ $88.70$ $95.23$
BLIP4CIR+Bi24’WACV BLIP BLIP $40.15$ $73.08$ $83.88$ $72.10$ $88.27$ $95.93$
SPIRIT24’TOMM CLIP(RN50) CLIP $40.23$ $75.10$ $84.16$ $73.74$ $89.60$ $95.93$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $46.22$ $78.17$ $87.64$ $70.92$ $87.69$ $94.68$
BLIP4CIR224’TMLR BLIP(ViT-B) BLIP $44.70$ $76.59$ $86.43$ $75.02$ $89.92$ $95.64$
CompoDiff(ST18M & FT)23’TMLR CLIP(ViT-L/14) CLIP $21.30$ $55.01$ $72.62$ $58.82$ $77.60$ $88.37$
CompoDiff(ST18M & FT)23’TMLR CLIP(ViT-G/14) CLIP $32.39$ $57.61$ $77.25$ $67.88$ $85.29$ $94.07$
SDQUR24’TCSVT BLIP2 $53.13$ $83.16$ $90.60$ $79.47$ $91.74$ $96.63$
NSFSE24’TMM ResNet152 BiGRU $20.70$ $52.50$ $67.96$ $44.20$ $65.53$ $78.50$
MANME24’TCSVT ResNet50 BiGRU $18.27$ $48.02$ $63.23$ $42.43$ $64.89$ $77.93$
SPRC24’ICLR BLIP2(ViT-L) $51.96$ $82.12$ $89.74$ $80.65$ $92.31$ $96.60$
Zero-Shot Learning-based CMR (ZSL-CMR)
PALAVAR22’ECCV CLIP(ViT-B/32) $16.62$ $43.49$ $58.51$
Pic2Word23’CVPR CLIP(ViT-L/14) $23.90$ $51.70$ $65.30$
SEARLE23’ICCV CLIP(ViT-B/16) $24.00$ $53.42$ $66.82$ $54.89$ $76.60$ $88.19$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $24.87$ $52.31$ $66.29$ $53.80$ $74.31$ $86.94$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $24.00$ $53.42$ $66.82$ $54.89$ $76.60$ $88.19$
MTCIR23’Arxiv CLIP(ViT-L/14) $25.52$ $54.58$ $67.59$ $55.64$ $77.54$ $89.47$
VQA4CIR(+CLIP4CIR)23’Arxiv $51.40$ $81.71$ $86.65$ $73.16$ $88.84$ $95.59$
VQA4CIR(+SPRC)23’Arxiv $54.00$ $84.23$ $91.85$ $82.07$ $93.45$ $97.08$
KEDs24’CVPR CLIP(ViT-L/14) $26.40$ $54.80$ $67.20$
PM24’Arxiv CLIP(ViT-L/14) $26.10$ $55.20$ $67.50$ $56.00$ $76.60$ $88.00$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $25.90$ $55.61$ $67.66$ $55.21$ $75.88$ $87.98$
Semi-Supervised Learning-based CMR (SSL-CMR)
MCL(OPT-6.7B)21’ICML $24.15$ $55.98$ $90.92$ $59.52$
MCL(Llama2-7B)21’ICML $26.22$ $56.84$ $91.35$ $61.45$
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $18.24$ $53.14$ $70.82$ $57.42$ $77.10$ $87.90$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $26.71$ $55.14$ $74.52$ $64.54$ $82.39$ $91.81$
TransAgg23’BMVC BLIP $38.10$ $68.42$ $93.51$
CASE24’AAAI BLIP(ViT)+BERT $48.00$ $79.11$ $87.25$ $75.88$ $90.58$ $96.00$
VISTA24’ACL $76.10$ $75.70$
HyCIR(CC3M+synthetic)24’Arxiv CLIP $25.08$ $53.49$ $67.03$ $53.83$ $75.06$ $87.18$
HyCIR(CC3M+synthetic)24’Arxiv BLIP $38.28$ $69.03$ $79.71$ $66.79$ $84.79$ $93.06$
VDG(Human+COCO’se)24’CVPR BLIP $49.37$ $78.12$ $85.52$ $76.68$ $90.46$ $96.05$
VDG(Human+NLVR2’se)24’CVPR BLIP $50.96$ $80.15$ $86.86$ $77.45$ $90.65$ $96.10$
PTG(SPRC+PTG)24’Arxiv $36.40$ $66.10$
TABLE VI: Comparison of Different Methods on CIRCO Dataset
Method Backbone CIRCO
Visual Textual mAP@1 mAP@10 mAP@25 mAP@50
Zero-Shot Learning-based CMR (ZSL-CMR)
PALAVAR22’ECCV CLIP(ViT-B/32) $4.61$ $5.32$ $6.33$ $6.80$
MTCIR23’Arxiv CLIP(ViT-L/14) $10.36$ $11.63$ $12.95$ $13.67$
Pic2Word23’CVPR CLIP(ViT-L/14) $8.72$ $9.51$ $10.64$ $11.29$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $10.18$ $11.03$ $12.72$ $13.67$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $11.68$ $12.73$ $14.33$ $15.12$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $15.05$ $16.32$ $18.06$ $19.05$
CIReVL24’ICLR CLIP(ViT-B/32) $14.94$ $15.42$ $17.00$ $17.82$
CIReVL24’ICLR CLIP(ViT-L/14) $18.57$ $19.01$ $20.89$ $21.80$
CIReVL24’ICLR CLIP(ViT-G/14) $26.77$ $27.59$ $29.96$ $31.03$
LinCIR24’CVPR CLIP(ViT-L/14) $12.59$ $13.58$ $15.00$ $15.85$
LinCIR24’CVPR CLIP(ViT-H/14) $17.60$ $18.52$ $20.46$ $21.39$
LinCIR24’CVPR CLIP(ViT-G/14) $19.71$ $21.01$ $23.13$ $24.18$
iSEARLE-XL-OTI24’Arxiv CLIP(ViT-L/14) $11.31$ $12.67$ $14.46$ $15.34$
iSEARLE-XL24’Arxiv CLIP(ViT-L/14) $12.50$ $13.61$ $15.36$ $16.25$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $11.26$ $12.11$ $13.63$ $14.37$
RTD(LinCIR)24’Arxiv CLIP(ViT-B/32) $8.94$ $9.35$ $10.57$ $11.21$
RTD(SEARLE)24’Arxiv CLIP(ViT-L/14) $16.53$ $17.89$ $19.77$ $20.68$
RTD(LinCIR)24’Arxiv CLIP(ViT-L/14) $17.11$ $18.11$ $20.06$ $21.01$
ISA(Sym)24’ICLR BLIP $9.67$ $10.32$ $11.26$ $11.61$
ISA(Asy)24’ICLR EfficientNet+BLIP $11.33$ $12.25$ $13.42$ $13.97$
Slerp24’ECCV CLIP(ViT-B/32) $6.35$ $7.11$ $8.12$ $8.75$
Slerp+TAT24’ECCV CLIP(ViT-B/32) $9.34$ $10.26$ $11.65$ $12.33$
Slerp+TAT24’ECCV CLIP(ViT-L/14) $18.46$ $19.41$ $21.43$ $22.41$
Slerp+TAT24’ECCV BLIP(ViT-L/16) $17.84$ $18.44$ $20.24$ $21.07$
LDRE24’SIGIR CLIP(ViT-B/32) $17.96$ $18.32$ $20.21$ $21.11$
LDRE24’SIGIR CLIP(ViT-L/14) $23.35$ $24.03$ $26.44$ $27.50$
LDRE24’SIGIR CLIP(ViT-G/14) $31.12$ $32.24$ $34.95$ $36.03$
HyCIR(Pic2Word-CLIP)24’Arxiv CLIP $14.12$ $15.02$ $16.72$ $17.56$
HyCIR(Pic2Word-BLIP)24’Arxiv BLIP $18.91$ $19.67$ $21.58$ $22.49$
Semi-Supervised Learning-based CMR (SSL-CMR)
MCL(OPT-6.7B)21’ICML $15.14$ $16.13$ $17.88$ $18.82$
MCL(Llama2-7B)21’ICML $17.67$ $18.86$ $20.80$ $21.68$
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $12.55$ $13.36$ $15.83$ $16.43$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $15.33$ $17.71$ $19.45$ $21.01$
TABLE VII: Performance Comparison of Different Methods on MIT-States and Birds-to-Words Datasets
Method Backbone MIT-States Birds-to-Words
Visual Textual R@1 R@5 R@10 Avg R@10 R@50
Supervised Learning-based CMR (SL-CMR)
TIRG18’CVPR ResNet18 LSTM $12.20$ $31.90$ $43.10$ $29.07$
LBF(small)20’CVPR Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
LBF(big)20’CVPR Faster-RCNN $14.72$ $35.3$ $46.56$ $32.19$
MAAF20’Arxiv ResNet50 LSTM $12.70$ $32.60$ $44.80$ $30.03$ $34.75$ $66.29$
TRACE21’AAAI ResNet50 GRU $19.56$ $45.24$
GSCMR21’TIP Faster-RCNN BiGRU $17.28$ $36.45$ $47.04$ $33.59$
ComposeAE21’WACV ResNet18 BERT $13.90$ $35.30$ $47.90$ $32.37$
MAN21’ICASSP ResNet18 BERT $13.90$ $35.30$ $47.90$ $31.53$
MAN21’ICASSP MobileNet BERT $15.60$ $36.70$ $47.70$ $33.33$
RTIC21’Arxiv ResNet50 LSTM $37.40$ $66.97$
HFCA21’ACMMM ResNet50 LSTM $14.30$ $35.36$ $47.12$ $32.26$
HCL21’MMAsia ResNet18 LSTM $15.22$ $35.95$ $46.71$ $32.63$
CRR22’ACMMM ResNet101 GRU $17.71$ $37.16$ $47.83$ $34.23$
TriArea22’Sci.Rep. ResNet18+TF LSTM $13.20$ $33.30$ $44.30$ $30.27$
SAC22’WACV ResNet50 GRU $19.56$ $45.24$
LIMN24’TPAMI ResNet50 LSTM $40.34$ $66.18$
6.3 Composed Video Retrieval
Composed Video Retrieval (CoVR) aims to retrieve specific videos from a large database by integrating both visual queries and modification texts, thus allowing for more sophisticated video searches. This approach addresses the limitations of traditional content-based video retrieval, which relies solely on visual features and often fails to fully capture user intent or context. The main goal of CoVR is to enhance the accuracy and relevance of video search results by leveraging multi-modal information. By combining a reference video with a textual description that specifies desired modifications or conditions, users can refine their search criteria to find videos that closely match their requirements. For instance, given a query image of a fountain in Barcelona, appending the text ”during show at night” would enable users to discover videos featuring that fountain during a nighttime show, offering a more precise and personalized search experience.
CoVR has significant application value across various domains, including e-commerce, fashion, internet video search, live event discovery, and sports video retrieval. In the e-commerce and fashion sectors, it helps customers quickly locate specific product videos showcasing particular attributes or usage scenarios, thereby enhancing the shopping experience. For online video platforms, this technology supports more effective content recommendation and management systems by accurately identifying and recommending videos that align with user interests and preferences.
6.3.1 Benchmark Datasets
WebVid-CoVR [91] is an extensive dataset created automatically from Web-scraped video-caption pairs, resulting in 1.6 million triplets. Typically, videos have a duration of 16.8 seconds, modification texts consist of 4.8 words, and each target video is linked to 12.7 triplets. WebVid-CoVR also features validation and test sets sourced from the WebVid10M corpus. The validation set contains 7,000 triplets, while the test set comprises 3,200 triplets that have been carefully curated to guarantee high quality. The dataset stands out for its scale and natural domain coverage, making it suitable for training robust models capable of handling real-world applications.
CIRR [42] and FashionIQ [82] are primarily focused on composed image retrieval but also serve as benchmarks for evaluating the zero-shot performance of CoVR models. These datasets provide valuable insights into how well CoVR methods can generalize to different types of data and tasks.
EgoCVR [92] is a meticulously curated benchmark designed for the task of Fine-Grained Composed Video Retrieval, utilizing an extensive egocentric video dataset. This dataset consists of 2,295 queries, specifically crafted to emphasize high-quality temporal video understanding. Each query and target clip is derived from the same long-form video, with textual modifications requiring subtle changes in the depicted actions, thus necessitating robust video comprehension capabilities for effective performance. EgoCVR’s construction involves collecting videos and corresponding narrations from the Ego4D dataset, focusing on diverse human-object interactions. To ensure the specificity and subtlety of action modifications, a rigorous manual annotation process was employed, contrasting with automated processes used in previous works. Consequently, this dataset highlights temporal modifications, with 1,811 samples (78.9%) centered on temporal events, as opposed to 484 samples (21.1%) involving object-centered changes. This focus significantly differs from the WebVid-CoVR-Test set, where 85% of samples concentrate on object modifications rather than temporal ones.
ICQ [93] is a pioneering benchmark specifically designed for the task of localizing events in videos using multimodal queries (MQs). This dataset is created to bridge the gap between traditional natural language queries (NLQs) and the more versatile MQs, which integrate images to represent semantic queries more flexibly. ICQ includes an evaluation dataset called ICQ-Highlight, featuring synthetic reference images and human-curated queries, offering a robust testbed for this novel task. The construction of ICQ involves utilizing a combination of reference images and refinement texts to form each query, enabling a more adaptable and nuanced application scenario. The dataset evaluates model performance across four distinct reference image styles to ensure comprehensive coverage of diverse scenarios. To accommodate existing video localization models for this new task, ICQ introduces three Multimodal Query Adaptation (MQA) methods and a Surrogate Fine-tuning on pseudo-MQs strategy, collectively known as SUIT. These approaches facilitate the integration of MQs into conventional NLQ-based backbone models.
6.3.2 Results and Analysis
6.4 Composed Remote Sensing Image Retrieval
Traditional remote sensing image retrieval (CRSIR) typically relies on either image content or text descriptions to find matching remote sensing images. However, this approach often falls short when precise identification of geographic features (e.g., shapes, colors, densities, targets) or specific time points is required. By integrating both image and text modalities, CRSIR allows users to submit both an example image and a textual description as query inputs. This significantly enhances the query’s descriptive power, enabling the system to interpret not only the visual content of the image but also the accompanying text, thereby refining the relevance of the search results. This method provides more accurate responses to users’ customized search needs and is applicable across various domains, including environmental monitoring, urban planning, and disaster response.
CRSIR has also proven to be highly beneficial in practical applications, offering both business and social advantages. It improves the efficiency and accuracy of image retrieval while enhancing the user experience by enabling more targeted decision-making through customized products and services. For geographic researchers, CRSIR accelerates research progress, optimizes resource utilization, and fosters interdisciplinary collaboration. In fields such as environmental monitoring, agricultural management, urban planning, and disaster response, CRSIR delivers exceptional data processing capabilities and substantial support for related research. In conclusion, CRSIR technology represents a significant advancement in remote sensing data applications, with both important academic implications and vast market potential.
6.4.1 Benchmark Datasets
PATTERNCOM [94] is a large-scale, high-resolution remote sensing image retrieval dataset consisting of 38 classes, with each class containing 800 images of 256×256 pixels. In PATTERNCOM, specific categories are selected to be described in the query image, and a corresponding query text defines attributes related to each class. For instance, the query image for the ”swimming pool” category is paired with text queries that describe attributes such as shape, with options like ”rectangular,” ”oval,” and ”kidney-shaped.” The dataset includes six attributes, each comprising up to four different classes. Each attribute can be associated with two to five values per class. The dataset contains a total of over 21,000 queries, with the number of positive queries per class ranging from 2 to 1345.
Airplane, Tennis, and WHIRT [95] are organized in terms of quintets, consisting of a reference RS image and its scene graph, a target RS image and its scene graph, and a pair of modifier sentences. Scene graphs capture object attributes and spatial relationships between pairs. Modifier sentences detail differences between reference and target images. Airplane contains 1600 images of airplanes and 3461 pairs of modifier sentences, and Tennis includes 1200 images of tennis courts and 1924 modifier sentence pairs, where both datasets are sourced from UCM [96], PatternNet [97], and NWPU-RESISC45 [98]. WHIRT comprises 4940 images from WHDLD [99] and 3344 manually generated ⟨reference images, and target image⟩ pairs.
6.4.2 Results and Analysis
6.5 Composed Person Retrieval
Composed Person Retrieval (CPR) represents an innovative approach to identifying specific individuals by leveraging both visual and textual information. Traditional methods, such as Image-based Person Retrieval (IPR) and Text-based Person Retrieval (TPR), often fall short in effectively utilizing both types of data, leading to a loss in accuracy. CPR aims to address this limitation by simultaneously employing image and text queries to enhance the retrieval process. This dual-modality approach not only increases the descriptive power of the query but also refines the relevance of search results, providing more accurate identification of target individuals. CPR is particularly useful in various applications, including social services and public security, where precise person identification is crucial.
The introduction of CPR has significant practical implications, offering substantial benefits for both business and social sectors. By improving the efficiency and accuracy of person retrieval, CPR enhances decision-making processes and user experiences through tailored solutions. For instance, in fields ranging from law enforcement to personalized marketing, CPR provides exceptional data processing capabilities and robust support for related research activities.
6.5.1 Benchmark Datasets
ITCPR [100] is a meticulously curated dataset designed for evaluating the performance of Zero-shot Composed Person Retrieval (ZS-CPR). It consists of finely annotated image-text pairs specifically created to address the challenges posed by CPR without requiring costly manual annotations. ITCPR leverages existing image-text datasets relevant to TPR tasks, ensuring a broad coverage of diverse scenarios. The dataset includes over 21,000 queries, each paired with corresponding images that depict the same identity under different outfits or conditions, thereby forming complete triplets (Iq, Tq, It). In ITCPR, reference images are combined with relative captions to form multimodal queries. These queries are designed to capture subtle variations in appearance, such as clothing changes or accessory additions, which are essential for distinguishing between individuals. Each triplet in the dataset comprises a reference image (Iq), a relative caption (Tq), and several target images (It) depicting the same individual in different contexts.
6.5.2 Results and Analysis
6.6 Composed Sketch-based Image Retrieval
The objective of Composed Sketch-Text Image Retrieval [101] is to combine sketch and textual information to enhance the accuracy and relevance of image retrieval. Specifically, it aims to leverage sketches to capture the shape and structure of objects, while textual descriptions supplement visual details such as color, material, and texture. This approach improves retrieval by integrating both coarse and fine-grained visual information. For instance, users may express their needs by sketching or describing an image, even in the absence of a specific picture.
The application value and practical use cases are diverse, including but not limited to the following areas: (1) In the design and creative industries, designers can use sketches to convey concepts and pair them with textual descriptions to aid in searching for inspiration or reference material within large databases, particularly in fields like fashion design and interior decoration. (2) In e-commerce and online shopping, consumers can provide hand-drawn sketches or textual descriptions of a product, allowing the system to recommend similar product images, thereby enhancing the shopping experience. (3) In healthcare, doctors can use this technology to identify rare diseases or medical devices by sketching the affected area or device shape and supplementing it with descriptive text, allowing for rapid location of relevant information to support diagnostic decisions. (4) In law enforcement, police can search for suspects based on eyewitness descriptions and suspect sketches, thereby improving investigative efficiency.
6.6.1 Benchmark Datasets
QMUL-ShoeV2 and QMUL-ChairV2 [102]: Both datasets emphasize detailed associations between sketch queries and corresponding images, with QMUL-ShoeV2 comprising 2000 sketches and 6730 photos, while QMUL-ChairV2 includes 400 sketches and 1800 photos.
Sketchy [103]: This dataset expands the scope by covering 125 categories, totaling 12,500 photos, each accompanied by at least five sketches.
FS-COCO [104] and SketchyCOCO [105]: FS-COCO contains 10,000 paired sketch-text-photo triplets, whereas SketchyCOCO offers 14,081 such triplets. The images and textual captions in these datasets originate from MS-COCO [106], providing a rich ground for studying multi-modal retrieval scenarios.
ImageNet-R(endition) [107]: Comprising 30,000 images across 200 ImageNet classes [108] and spanning 16 domains, this dataset is particularly useful for examining domain-specific attributes and their transferability in image retrieval contexts.
6.6.2 Results and Analysis
6.7 Interactive/Conversational Retrieval
Interactive/Conversational Retrieval (ICR) represents an advanced approach to image retrieval that leverages natural language interactions between users and systems to refine search outcomes progressively [84, 109, 110, 111, 112, 113, 114, 115, 116, 117]. Unlike traditional methods relying solely on images or predefined textual attributes, ICR integrates user feedback through conversational interfaces, enhancing the accuracy and relevance of search results. This method enables users to provide iterative feedback in natural language, refining queries dynamically until they locate the desired image or item. The primary objective of ICR is to facilitate more intuitive, precise, and personalized searches by incorporating both visual and semantic information effectively.
ICR has significant applications across various domains, including e-commerce, fashion, and social media, where users often seek specific items with complex characteristics such as style, color, and material. By enabling multi-turn interactions, ICR improves user experience and satisfaction by allowing for more detailed and nuanced queries. For instance, in fashion retail, customers can start with a reference image of a clothing item and iteratively modify their query (”long-sleeve version”, ”in blue color”) to find exactly what they are looking for. This not only enhances the shopping experience but also supports retailers in providing accurate product recommendations based on user preferences, thereby boosting customer engagement and sales.
6.7.1 Benchmark Datasets
Several benchmark datasets have been developed to evaluate and advance the capabilities of Interactive/Conversational Retrieval systems. These include:
Multi-turn FashionIQ [111], which extends the original FashionIQ [82] dataset into a multi-turn setting. It includes 11,505 sessions across three clothing types, structured into transactions of 2-turns, 3-turns, and 4-turns. Each session’s data is represented as a pair (In, Un), where In and Un denote the query image and the feedback text for each turn, respectively. This dataset is particularly useful for evaluating systems that require iterative refinement of search queries over multiple rounds of interaction.
Multi-turn Shoes: The initial Shoes dataset [118] encompasses a collection of images featuring women’s footwear, spanning 10 distinct categories, which were sourced online and automatically annotated with various attributes. To enhance the applicability of these images for single-turn, feedback-oriented image retrieval tasks, [118] introduced supplementary natural language descriptions, creating about 10k training pairs and 4.6k test queries. Within the scope of this work, Pal et al. [113] extended this dataset to accommodate multi-turn interactions, linking multiple single-turn interactions by matching the target image from one session to the query image of another. This extension facilitates research into multi-turn retrieval scenarios by concatenating several single-turn transactions and maintaining consistency with Multi-turn FashionIQ, thus offering valuable insights into memory retention and feedback handling across multiple turns.
Interactive Retrieval [110] is meticulously designed to support multi-turn interactive retrieval tasks involving complex scenes. Originating from the Visual Genome [krishna2017visual], this dataset includes detailed captions for various regions within each image, making it ideal for scenarios where users refine queries based on feedback. Comprising 8,960 training samples, approximately 10% of the entire dataset, it evaluates model performance under limited labeled data conditions. For rigorous assessment, a subset featuring over 18 nouns per caption, totaling 7,000 samples, emphasizes challenging cases with potential confounders and spurious correlations. A user simulator mimics real human interactions by providing image-click and text-click feedback on candidate images. Validated through 2,500 user sessions conducted by six annotators, the simulator accurately predicts actual user clicks in 85% and 71.4% of sessions, respectively.
7 Conclusion
In this paper, we provided a comprehensive review of composed multi-modal retrieval (CMR), an emerging research field that integrates visual and textual information to achieve more flexible and precise content-based retrieval. We discussed the evolution from unimodal to cross-modal and then to composed multi-modal retrieval, highlighting its significance and wide-ranging applications in e-commerce, social media, and beyond. Furthermore, we explored key research directions, including supervised, zero-shot, and semi-supervised learning-based CMR, analyzing their methodologies, challenges, and advancements.
Despite the remarkable progress, CMR still faces several open challenges, such as improving retrieval accuracy under zero-shot settings, handling large-scale datasets efficiently, and ensuring better generalization across different domains. Moreover, future research should focus on leveraging advancements in foundation models, generative AI, and self-supervised learning to enhance retrieval performance. We believe that continued exploration of CMR will play a pivotal role in shaping next-generation retrieval systems, enabling more intuitive and effective interactions between humans and vast multimodal information spaces.